### v1

# Neural variational inference for cutting feedback during uncertainty propagation

Abstract
In many scientific applications, uncertainty of estimates from an earlier (upstream) analysis needs to be propagated in subsequent (downstream) analysis, without feedback. Cutting feedback methods, also termed cut-Bayes, achieve this by constructing a cut-posterior distribution that prevents backward information flow, facilitating unidirectional uncertainty propagation. However, cutting feedback using sampling methods like nested MCMC is computationally challenging. Recently variational inference (VI) methods have been developed to approximate the cut-posterior. However, they need two variational approximations and require access to the upstream data and model. In this manuscript we propose, NeVI-Cut, a provably accurate and modular neural network-based variational inference method for cutting feedback. We directly utilize samples from the upstream analysis without requiring access to the upstream data or model. This simultaneously preserves modularity of analysis and reduces approximation errors by avoiding a variational approximation for the upstream model. Conditioning on each upstream sample, we then use normalizing flows (neural network-based generative models) to specify the conditional variational family for the downstream parameters. We estimate the conditional cut-posterior by optimizing the Monte Carlo average of the variational loss over all the upstream samples. We provide theoretical guarantees on the NeVI-Cut estimate to approximate any cut-posterior. Our results are in a fixed-data regime. We do not assume any asymptotic paradigm or even sample size growth, and go beyond just establishing universal expressiveness of normalizing flows. We provide convergence rates of the actual variational solution and quantify how richness of the neural architecture and the complexity of the target cut-posterior dictate the approximation quality. In the process, we establish new results on uniform Kullback-Leibler approximation rates of conditional distributions by common flow classes like Unconstrained Monotonic Neural Network and neural Rational Quadratic Splines. A triply stochastic scalable algorithm is devised to implement the method. Simulation studies and two real-world analyses illustrate how NeVI-Cut achieves significant computational gains over traditional cutting feedback methods and is considerably more accurate than parametric variational cut approaches.
Keywords: Bayesian analysis, Cutting Feedback; Modular Inference; Variational Inference; Normalizing Flows; Neural Networks, Universal approximation rates.
1 Introduction
Across many scientific fields, estimates of quantities from an upstream analysis are prestored to be reused in many downstream analyses. Often these upstream parameter estimates are quantified with uncertainty (Monte Carlo samples, ensemble members, or Bayesian posterior samples) and any use of these parameters in any downstream analysis mandates proper propagation of uncertainty. For example, downstream climate impacts studies (e.g., Butler et al.,, 2022; Park et al.,, 2023; Rao et al.,, 2024) commonly use ensemble outputs from climate models such as CMIP6 (Thrasher et al.,, 2022) as probabilistic projection samples, rather than using single climate trajectories, so that the spread from the climate model propagates into sectoral outcomes. Similarly, in environmental health and air quality research, widely used data products explicitly quantify uncertainty. The Global Burden of Disease project releases fine particulate matter (PM2.5) exposure and risk estimates as $1000$ posterior draws (Institute for Health Metrics and Evaluation,, 2024) that are used downstream in attributable-burden calculations (Murray et al.,, 2020). Other widely used high-resolution exposure surface data products also provide estimates with uncertainty (PM2.5 and NO2 estimates developed in Di et al., 2019b, ; Di et al., 2019a, ). Importance of incorporating this uncertainty in downstream health outcomes analysis has been routinely emphasized in the related literature on exposure measurement error (Peng and Bell,, 2010; Chang et al.,, 2011; Spiegelman,, 2016; Comess et al.,, 2024). Other examples of uncertainty quantified prestored outputs include posterior distributions of cosmological parameters, termed as Planck MCMC chains, posterior samples of confusion matrices of cause-of-death classifiers (Pramanik et al., 2025a, ), and probabilistic word-embeddings used in large language models (LLM) like word2Gauss and Bayesian skip-gram (Vilnis and McCallum,, 2014; Bražinskas et al.,, 2018).
In many of the aforementioned applications, when propagating uncertainty to a downstream Bayesian analysis, it is desirable not to have feedback, which is the statistical phenomenon of the downstream data informing an updated distribution for the upstream parameters. This is because the upstream quantities are often more fundamental and causally precedent (e.g., air pollution concentrations or climate data) than downstream ones (e.g., individual-level health outcomes). Additionally, downstream models are often misspecified for the upstream quantities. For example, in downstream health association studies, upstream estimates of area-level air pollution outputs are often used only as proxies for the true unobserved individual-level exposures (Zeger et al.,, 2000). Similarly, confusion matrix estimates of classifiers trained on specific populations are often used in downstream calibration of different target populations (Pramanik et al., 2025a, ). Hence, it is undesirable to update these upstream parameter distributions based on the downstream data and model that are subject to model misspecifications, dataset shifts, and biases. Furthermore, many of these upstream quantities (e.g., climate outputs, air pollution estimates) are used in many different downstream analyses, and it is impractical to expect revised upstream posteriors to be published after every downstream study conducted by different groups. In such settings, when conducting a downstream Bayesian analysis using uncertainty-quantified upstream quantities, one should prevent feedback and preserve the original upstream posterior while coherently propagating its uncertainty to the downstream quantities.
Feedback occurs in sequential Bayesian analysis if the upstream parameter distribution is used as a prior for that parameter in the downstream analysis (e.g., the analysis in Pramanik et al., 2025b, ) or if the upstream samples are used downstream using recursive Bayes methods (see e.g., Taylor et al.,, 2025, for a recent work). Both of these approaches approximate a full Bayesian analysis using both upstream and downstream data, which allows information sharing in all directions to infer all parameters jointly from all data, thereby resulting in feedback. The idea of explicitly stopping downstream to upstream feedback is known as Cutting feedback (Liu et al.,, 2009; Plummer,, 2015; Jacob et al.,, 2017). Instead of trying to estimate the conventional Bayesian posterior, cutting feedback methods target the cut-posterior, a modification that yields a valid joint distribution over all parameters while preserving the distribution of the upstream parameters from the first analysis.
There is now a large and rapidly growing literature on cutting feedback and related methods because of their broad applicability. Cutting feedback methods has been applied in various areas, including pharmacology (Lunn et al.,, 2009), economics (Murphy and Topel,, 2002), environmental health (Peng and Bell,, 2010, their two-stage Bayesian model is a version of nested MCMC used in cut models, although they didn’t make the explicit connection), and causal inference (Zigler,, 2016). Moss and Rousseau, (2024) use cut-posteriors in a hidden Markov model, preventing feedback from the emission distribution to transition parameter estimation. Smith et al., (2025) propose cutting feedback methods for misspecified copula models, improving accuracy compared to conventional Bayesian posteriors. Carmona and Nicholls, (2020, 2022) introduced semi-modular inference, that interpolates posteriors between Bayes and cut-Bayes, allowing partial feedback regulated by a hyperparameter for greater flexibility. More generally, cut-Bayes can be viewed as a case of modular Bayesian inference involving several data-model pairs where certain directions of information flow are unwarranted. Liu and Goudie, (2025) present a general framework for this, formally defining modules via Markovian relationships or directed acyclic graphs, and developing methods for general modular inference. We refer to Nott et al., (2023) for an overview of recent developments in Bayesian inference for misspecified models, including cutting feedback methods.
A major thread of developments on cutting feedback methods has been computational. The cut function in the OpenBUGS package modifies the MCMC conditional updates to restrict feedback between modules; however, Plummer, (2015) showed that this algorithm does not converge to the target cut-posterior. In general, designing a Markov Chain Monte Carlo (MCMC) that has the cut-posterior as the invariant distribution can be challenging. A principled alternative is the multiple imputation-based nested MCMC. This involves first running an MCMC for the upstream analysis to obtain a set of posterior samples, and then for each of these samples, running an MCMC for the downstream analysis, and then combining over all the MCMC posteriors. The repeated MCMC runs for the downstream analysis significantly increase computational cost. Plummer, (2015) proposed a tempered cut algorithm for approximating the cut-posterior. Jacob et al., (2020) developed an unbiased MCMC estimator using coupled chains and demonstrated the use for approximating the cut-posterior. Liu and Goudie, (2022) introduced a stochastic approximation adaptive MCMC method to approximate the cut-posterior, improving convergence and stability compared with the openBUGS algorithm and tempered cut. Pompe and Jacob, (2021) designed a posterior bootstrap method for modular inference, providing valid uncertainty quantification. Chakraborty et al., (2023) introduce a likelihood-free approach using Gaussian mixtures to approximate cut and semi-modular posteriors when the likelihood is intractable. Hutchings et al., (2025) proposed the Emulating the Conditional Posterior (ECP) algorithm, which uses emulation to increase the number of imputations, improving sampling efficiency for cut-posteriors. Mathews et al., (2025) designed a sequential Monte Carlo method to compute cut-posteriors.
Variational inference (VI), an optimization-based alternative to sampling methods for estimating distributions, has been explored in the context of cutting feedback to address the computational challenges. Yu et al., (2023) proved a fundamental result, showing that the cut-posterior can be viewed as a constrained variational optimization problem. They approximate the cut-posterior using fixed-form parametric (e.g., Gaussian) variational families. Variational inference was also used in Smith et al., (2025) for cutting feedback in misspecified copula models. In semi-modular inference (SMI), Carmona and Nicholls, (2022) used variational inference with normalizing flows, a class of neural network-based distributions, to approximate SMI posteriors flexibly. Battaglia and Nicholls, (2024) extended the flow-based approach to amortized inference over prior hyperparameters, fitting a conditional normalizing flow conditioning on the hyper-parameter values.
The existing variational inference approaches for cut-Bayes require access to the upstream raw data and model. These may not be available in many of the applications discussed above, where only the uncertainty quantified outputs are published (e.g., climate-model outputs, air pollution estimates, cosmological parameters, probabilistic word-embeddings). Sometimes the raw data may not be shared because of privacy concerns related to human-subjects research (e.g., uncertainty quantified estimates of cause-of-death classifiers published in Pramanik et al., 2025a, , are based on individual-level cause-of-death data that are not publicly available). Even if the upstream raw data and model are available, there is no need for the first variational approximation, as the cut-posterior for the upstream quantities is the same as their Bayes posterior from the upstream analysis. Also, many variational cut-Bayes methods rely on parametric assumptions for the variational family, such as the mean-field approximation that approximates marginal posteriors adequately but often fails to capture the complex conditional distributions. This is problematic, particularly when the true conditional cut-posterior deviates substantially from the chosen parametric form.
In this manuscript, we propose a modular neural network-based variational inference approach for cut-Bayes that mitigates these limitations. Our contributions are twofold. First, we directly utilize samples of parameter distributions from the upstream analysis, and construct a single loss function that integrates all samples. When optimized, this estimates the variational conditional cut-posterior of the downstream parameters. This eliminates the need to access the upstream raw data and model, thereby enabling modularity of cut-Bayes analysis and protecting data privacy. Even if the upstream data and model are available, we simply use a single MCMC run to obtain upstream samples and use these in the loss for the downstream variational analysis. Thus, unlike the current approaches, we do not need a variational approximation for the upstream posterior thereby reducing additional approximation error. We use conditional normalizing flows, neural network-based conditional distribution families. to estimate the conditional cut-posterior of the downstream parameters. Our proposed method, Neural Variational Inference for Cut-Bayes (NeVI-Cut), thus leverages the expressive power of neural networks (Hornik,, 1991) and normalizing flows (Rezende and Mohamed,, 2015) enabling better approximation of complex, multi-modal, or skewed posteriors. Computationally, NeVI-Cut demonstrates significant improvement over methods like the multiple imputation-based nested MCMC for cutting feedback, which can also work only with upstream samples.
A second contribution of this manuscript is developing a comprehensive theory for NeVI-Cut and variational cut-Bayes. Prior theoretical work on variational cut-Bayes is limited. Smith et al., (2025) prove consistency and asymptotic normality of the parameters of the Gaussian variational class for copula models, but their guarantees are asymptotic, appealing to Bernstein-von Mises type arguments that the true cut-posterior is approximately Gaussian asymptotically, thereby justifying the Gaussian variational family. Such results are reassuring yet do not address the quantity of real interest in applications, namely the quality of approximation to the exact cut-posterior conditional on the actually observed dataset, rather than a hypothetical large sample limit. On the other hand, Battaglia and Nicholls, (2024) provides a universal approximation result for conditional normalizing flows in terms of convergence in distribution. However, that is not sufficient to offer guarantees about the variational solutions using these flows, which minimize the Kullback-Leibler (KL) loss.
Our theory is in the ‘fixed-data’ or ‘fixed-target’ regime, not asymptotic and not assuming any sample size growth. This is the more challenging regime of theory when studying the approximation quality of a variational solution to the target posterior. In an asymptotic paradigm, Bernstein-von Mises implies approximate Gaussian posteriors, so even simple Gaussian VI can be asymptotically adequate. That safety net is absent in the fixed target setting, where posteriors can have complicated shapes requiring sufficiently complex variational classes like normalizing flows. We keep the observed data fixed, and provide guarantees on the approximation quality of the variational estimate obtained from NeVI-Cut with respect to the cut-posterior conditional on the observed data. Our results explicitly quantify how the approximation rates depend on both the complexity of the target cut-posterior as well as the richness (number of parameters) of the neural architecture used to design the conditional flows. In the process, we establish, to our knowledge, novel results on universal and uniform approximation rates of conditional flows in the Kullback-Leibler divergence (KLD). Our results cover common conditional flow classes like Unconstrained Monotonic Neural Network flows (UMNN, Wehenkel and Louppe,, 2019) and neural Rational Quadratic Splines (RQS, Durkan et al.,, 2019). As the existing literature on flows and conditional flows (Huang et al.,, 2018; Papamakarios et al.,, 2021) has mostly focused on their representational capabilities, our results are of independent importance to study variational solutions using flows.
The remainder of this paper is organized as follows: In Section 2, we illustrate the idea of cutting feedback. We then give background on variational inference and normalizing flows. In Section 3 we present our NeVI-Cut algorithm. Theoretical results are presented in Section 4. Section 5 details simulation experiments to illustrate the behavior of our model compared with the multiple imputation-based nested MCMC, the full Bayes approach, and parametric VI methods for cut-Bayes. Section 6 presents two real-world analyses using cutting feedback methods: a commonly used HPV data analysis, and the child and neonatal mortality data from the Countrywide Mortality Surveillance for Action (COMSA) Program in Mozambique. Section 7 concludes the paper with a discussion. Code to implement our proposed method is available on GitHub.
2 Background
2.1 Cutting Feedback Approaches
We illustrate the concept of cutting feedback from downstream to upstream in Figure 1. Black arrows indicate parameters specifying a model, blue dotted arrows indicate the direction of information flow from a data source to parameters. Let $D_{1}$ denote the upstream data which informs the upstream quantity $\eta$. Here $\eta$ can be a set of parameters that enters the model for $D_{1}$ (in which case there would be a black arrow going from $\eta$ to $D_{1}$ and which is the standard cut model schematic used previously, e.g., in Plummer, (2015)). However, $\eta$ need not be a parameter in the model for $D_{1}$, it can be simply quantities that are informed by $D_{1}$ (e.g., derived quantities like predicted variables based on $D_{1}$ and its model). Our approach is agnostic to this. Hence, we do not specify a black arrow from $\eta$ to $D_{1}$ and simply show the eventual direction of information flow in the upstream analysis, that $\eta$ is informed by $D_{1}$. The downstream data is denoted by $D_{2}$ whose model involves both $\eta$ and some new parameters $\theta$ (hence the black arrows from $\eta$ and $\theta$ to $D_{2}$ in the left figure).
Figure 1: Paradigm of the cutting feedback from downstream to upstream.
The right figure shows the final directions of information flow. If both datasets were analyzed jointly, the Bayesian posterior distribution can be written as:
$p(\theta,\eta\,|\,D_{1},D_{2})\,\,\propto\,\,p(\theta\,|\,\eta,D_{1},D_{2})\,p(\eta\,|\,D_{1},D_{2})=p(\theta\,|\,\eta,D_{2})\,p(\eta\,|\,D_{1},D_{2}).$ (1)
As we know in a Bayesian analysis, the final posterior of $\eta$ is informed by both $D_{1}$ and $D_{2}$ (hence the blue arrows from both $D_{1}$ and $D_{2}$ to $\eta$) and is thus different from its posterior $p(\eta\,|\,D_{1})$ after the upstream analysis. As discussed in the introduction, this is undesirable in many situations. For example, if the second module is misspecified (which is the case in many of the applications we highlighted), inference on $\eta$ could be negatively affected from this backward information flow from $D_{2}$ to $\eta$.
Cutting feedback from $D_{2}$ to $\eta$ (as indicated by the red dotted line) prevents this. In other words, instead of considering the Bayes posterior $p(\eta\,|\,D_{1},D_{2})$, we modify it to propagate the uncertainty in $\eta$ into the inference of $\theta$ from $D_{2}$ without feedback from $D_{2}$ to $\eta$. The marginal cut-posterior for $\eta$ is thus defined $p_{\text{cut}}(\eta\,|\,D_{1},D_{2})=p(\eta\,|\,D_{1})$, i.e., the posterior remains unchanged from the upstream analysis. We can then get the following joint cut-posterior:
$p_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})$ $=\,p(\theta\,|\,\eta,D_{1},D_{2})\,p_{\text{cut}}(\eta\,|\,D_{1},D_{2})$ (2)
$=p(\theta\,|\,\eta,D_{2})\,p(\eta\,|\,D_{1}).$
Here, $p(\theta\,|\,\eta,D_{1},D_{2})=p(\theta\,|\,\eta,D_{2})$, as in (1), because the likelihood of $D_{1}$ does not depend of $\theta$ (hence no arrows between them), and as mentioned above, $p_{\text{cut}}(\eta\,|\,D_{1},D_{2})=p(\eta\,|\,D_{1})$ preserves the posterior of $\eta$ from upstream analysis and cuts the feedback from downstream. The marginal cut-posterior for $\theta$ is then given by
$p_{\text{cut}}(\theta\,|\,D_{1},D_{2})$ $=\int p(\theta\,|\,\eta,D_{2})\,p(\eta\,|\,D_{1})d\eta.$ (3)
Hence, to sample from the joint cut-posterior of $(\eta,\theta)$ or for the marginal cut-posterior of $\theta$, one can generate draws from $p(\theta\,|\,\eta,D_{2})$ for values of $\eta$ from $p(\eta\,|\,D_{1})$. This propagates the uncertainty of $\eta$ from the upstream analysis into the marginal cut-posterior for $\theta$. However, $\theta$ is typically multidimensional and direct sampling from $p(\theta\,|\,\eta,D_{2})$ is generally not feasible. Hence, for each draw of $\eta$ from $p(\eta\,|\,D_{1})$ (obtained from an MCMC run), a second nested MCMC is run to obtain samples of $\theta\,|\,\eta,D_{2}$. This nested MCMC scheme to sample from the cut-posterior is referred to as multiple imputation (MI; Plummer,, 2015) and is very computationally intensive as it requires MCMC runs for each sample for $\eta$ from the upstream analysis.
2.2 Variational Inference
Variational inference (VI) provides a practical alternative by approximating posterior distributions via optimization (Blei et al.,, 2017). Given data $D$ with model specified by some parameters $\phi$, VI seeks a distribution $q(\phi)$ within a variational family $\mathcal{Q}$ that minimizes the KLD to a posterior distribution $p(\phi\,|\,D)$, to get the optimal solution:
${q}^{*}(\phi)=\underset{q(\phi)\in\mathcal{Q}}{\arg\min}\,\,\mathrm{KL}\,\!({q}(\phi)\parallel p(\phi\,|\,D)).$ (4)
If no restriction is placed on the family $\mathcal{Q}$ of distributions, then the solution $q^{*}(\phi)$ is exactly the true posterior distribution $p(\phi\,|\,D)$. In practice, one places some parametric assumptions on the class $\mathcal{Q}$ to obtain an optimizer that approximates the true posterior distribution.
A fundamental contribution in cut-Bayes or cutting feedback methods that enabled the use of variational inference for this problem is the result (Lemma 4.1) of Yu et al., (2023). Considering the setup of Section 2.1 with $\phi=(\theta,\eta)$, the result proved that when $\mathcal{Q}$ is restricted to all distributions $q(\theta,\eta)$ such that $q(\eta)=p(\eta\,|\,D_{1})$, the optimization in (4) yields the joint cut-posterior (2) as the solution. Formally,
$p_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})={q}^{*}(\theta,\eta)=\underset{q(\theta,\eta)\in\mathcal{Q}:q(\eta)=p(\eta\,|\,D_{1})}{\arg\min}\,\,\mathrm{KL}\,\!({q}(\theta,\eta)\parallel p(\theta,\eta\,|\,D_{1},D_{2})).$ (5)
Thus, the joint cut-posterior can be characterized as the variational solution, minimizing KLD to the joint full Bayes posterior under the constraint of no feedback from $D_{2}$ to $\eta$. In their implementation, they used two variational approximations, decomposing $q(\theta,\eta)=q(\theta\,|\,\eta)\,q(\eta)$, using parametric families for both, i.e., $q(\theta\,|\,\eta)=q_{\lambda_{2}}(\theta\,|\,\eta)$ and $q(\eta)=q_{\lambda_{1}}(\eta)$. To estimate $q^{*}(\eta)$, the KLD to $p(\eta\,|\,D_{1})$ was minimized with respect to $\lambda_{1}$ to obtain $\hat{q}^{*}(\eta)=q_{\hat{\lambda}_{1}}(\eta)$. Then $q_{\hat{\lambda}_{1}}(\eta)$ is plugged into the second optimization (5) to estimate the minimizing parameters $\hat{\lambda}_{2}$. The final cut-posterior estimate is then given by $q_{\hat{\lambda}_{2}}(\theta\,|\,\eta)\,q_{\hat{\lambda}_{1}}(\eta)$. Similar ideas were used in Smith et al., (2025) in the context of copula models.
2.3 Normalizing Flows
The effectiveness of variational inference depends on the choice of variational family. Conventional choices such as mean-field approximations, i.e., assuming $q(\phi)=\prod_{i}q_{i}(\phi_{i})$, often fail to capture complex posterior structures. Normalizing flows provide a flexible alternative by constructing complex densities through a sequence of invertible and differentiable transformations of a simple base distribution, and have been widely explored in variational inference (Rezende and Mohamed,, 2015).
The basic idea of normalizing flows is as follows. If $F$ is the cumulative distribution function (cdf) of a scalar $\theta$, then there always exists a function $f$ such that $\theta\overset{d}{=}f(Z)$, where $Z$ is some base distribution. The function $f$ is explicitly obtained from the inverse cdf transformation, e.g., if $Z\sim N(0,1)$, $f(Z)=F^{-1}(\Phi(Z))$ where $\Phi$ is the standard Normal cdf. Hence, the task of approximating the distribution of $\theta$ becomes equivalent to approximating the function $f$. As neural networks are universal approximators of many function classes, it thus suffices to model $f$ using a suitable class of neural networks $g_{\lambda}$ and estimate the neural parameters (weights and biases) $\lambda$ using the variational (KL) loss. Normalizing flows build on this core idea, but the specific choices to model $f$ vary widely. We will discuss some common flow classes when introducing our method and studying its theoretical properties. A more comprehensive review of normalizing flows and their extensions can be found in (Kobyzev et al.,, 2020; Papamakarios et al.,, 2021).
3 Neural Variational Inference for Cut-Bayes
In this section, we propose a variational inference method for cutting feedback that $(a)$ can directly use samples of $\eta$ from the upstream posterior $p(\eta\,|\,D_{1})$ without requiring access to the raw data and model; and $(b)$ uses normalizing flows to non-parametrically model the conditional cut-posterior of $\theta$ given $\eta$.
Recall from (5) that the cut-posterior is the solution to the variational optimization with the variational family $q(\theta,\eta)$ constrained to have $q(\eta)=p(\eta\,|\,D_{1})$ to prevent feedback. As $q(\eta)$ is fixed, we make two observations. First, the optimization in (5) is essentially only over the class of conditional distributions $q(\theta\,|\,\eta)$. Also, under this constraint of $q(\eta)=p(\eta\,|\,D_{1})$, the KL loss function in (5) can be written as $\mathrm{KL}\,\!({q}(\theta,\eta)\parallel p(\theta,\eta\,|\,D_{1},D_{2}))=\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\big[\mathrm{KL}\,\!(q(\theta\,|\,\eta)\parallel p(\theta\,|\,\eta,D_{2}))\big]$.
Hence, we can express the joint cut-posterior in terms of the conditional variational solution as
$p_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})$ $=q^{*}_{\text{cut}}(\theta\,|\,\eta)\,p(\eta\,|\,D_{1}),\mbox{ where }$ (6)
$q^{*}_{\text{cut}}(\theta\,|\,\eta)$ $=\underset{q(\theta\,|\,\eta)}{\arg\min}\;\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\left[\mathrm{KL}\,\!(q(\theta\,|\,\eta)\parallel p(\theta\,|\,\eta,D_{2}))\right].$
Thus, finding the joint cut-posterior essentially reduces to estimating the collection of conditional distributions $\{q(\theta\,|\,\eta)\}_{\eta}$. No approximations to $p(\eta\,|\,D_{1})$ is needed. In practice, we will typically have samples $\eta_{1},\ldots,\eta_{N}$ from $p(\eta\,|\,D_{1})$. As long as the samples are iid or from an ergodic MCMC, we can approximate the expectation term in (6) with the sample average, leading to the estimate
$\hat{p}_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})$ $=\hat{q}_{\text{cut}}(\theta\,|\,\eta)\,p(\eta\,|\,D_{1}),\mbox{ where }$ (7)
$\hat{q}_{\text{cut}}(\theta\,|\,\eta)$ $=\underset{q(\theta\,|\,\eta)}{\arg\min}\;\frac{1}{N}\sum_{i=1}^{N}\left[\mathrm{KL}\,\!(q(\theta\,|\,\eta_{i})\parallel p(\theta\,|\,\eta_{i},D_{2}))\right]$
$=\underset{q(\theta\,|\,\eta)}{\arg\min}\;\frac{1}{N}\sum_{i=1}^{N}\left[\mathrm{KL}\,\!(q(\theta\,|\,\eta_{i})\parallel p(D_{2}\,|\,\theta,\eta_{i})\,p(\theta\,|\,\eta_{i})\right].$
Equations (6) and (7) illustrate how the joint cut-posterior can be estimated without access to the upstream data $D_{1}$ and the model used in upstream analysis, as long as the samples $\{\eta_{i}\}$ are available. Even if $D_{1}$ and its model are available, there is no need to approximate $p(\eta\,|\,D_{1})$ using the additional variational inference step, as done in Yu et al., (2023); Smith et al., (2025). One can run a standard MCMC, to estimate $p(\eta\,|\,D_{1})$ and use the posterior samples $\{\eta_{i}\}$ in (7). Once the downstream analysis model likelihood $p(D_{2}\,|\,\theta,\eta)$ and the prior $p(\theta\,|\,\eta)$ are chosen, the only unknown in (7) is the class of conditional distributions $q(\theta\,|\,\eta)$. In the next section, we propose using normalizing flows that can approximate any conditional family, leading to a provably accurate estimate of the joint cut-posterior.
3.1 Normalizing Flows for Conditional Distribution
We propose modeling $q(\theta\,|\,\eta)$ using conditional normalizing flows, which approximate complex distributions using transformations of simple base distributions. Specifically, if $\theta\in\mathbb{R}^{d}$, we consider a class of conditional distributions $q_{\vartheta}(\theta\,|\,\eta)$ that corresponds to the law $\theta\,|\,\eta\sim T_{\vartheta}(\eta,Z)$. Here $T_{\vartheta}(\eta,Z)$ is a class of neural network-based functions (choices discussed later) with parameters $\vartheta$ taking values in some space $\Theta$, and $Z\in\mathbb{R}^{d}$ is a latent variable of the same dimension as $\theta$ and with a base distribution $p(Z)$ that is easy to sample from. A common choice is the standard normal distribution, e.g., $Z\sim N(0,I_{d})$. However, if the true posterior is heavy-tailed, then Cauchy or Student’s $t$ may be a better base. See Section 4 for theoretical insights on the choice of the base.
Using the classical reparametrization trick of variational inference (Kingma et al.,, 2013), the Kullback Leibler objective in (7), can be expressed in terms of $p(Z)$ as
$\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\,\mathbb{E}_{Z\sim p(Z)}\big[\log q_{\vartheta}\!\left(T_{\vartheta}(\eta,Z)\mid\eta\right)-\log p\!\left(T_{\vartheta}(\eta,Z)\mid\eta,D_{2}\right)\big].$ (8)
Writing $J_{z}T_{\vartheta}(\eta,z)$ for the Jacobian of $T_{\vartheta}$ with respect to $z$, and using change of variables (see Lemma 2 for a formal statement and proof), we have the identity
$\log q_{\vartheta}\!\left(T_{\vartheta}(\eta,z)\mid\eta\right)=\log p(z)-\log\!\left|\det J_{z}T_{\vartheta}(\eta,z)\right|.$
Plugging this in (8), we can write the loss as
$\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\,\mathbb{E}_{Z\sim p(Z)}\big[\log p(Z)-\log|\det J_{z}T_{\vartheta}(\eta,Z)|-\log p\!\left(T_{\vartheta}(\eta,Z)\mid\eta,D_{2}\right)\big].$
Finally, noting that $\log p(Z)$ does not involve any terms of $\vartheta$ and $\log p\!\left(T_{\vartheta}(\eta,Z)\mid\eta,D_{2}\right)=\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta,Z),\eta\right)+\log p\!\left(T_{\vartheta}(\eta,Z)\,|\,\eta\right)+$ terms free of $\vartheta$, minimizing (8) is equivalent to maximizing the average conditional evidence lower bound (ELBO)
$\mathcal{L}(\vartheta)=\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\,\mathbb{E}_{Z\sim p(Z)}\big[\log|\det J_{z}T_{\vartheta}(\eta,Z)|+\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta,Z),\eta\right)+\log p\!\left(T_{\vartheta}(\eta,Z)\,|\,\eta\right)\big].$
Once again, we can replace the expectation under $p(\eta\,|\,D_{1})$ with average over the samples $\{\eta_{i}\}$, leading to the objective
$\hat{\mathcal{L}}(\vartheta)=\frac{1}{N}\sum_{i=1}^{N}\,\mathbb{E}_{Z\sim p(Z)}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z)|+\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta_{i},Z),\eta_{i}\right)+\log p\!\left(T_{\vartheta}(\eta_{i},Z)\,|\,\eta_{i}\right)\big].$ (9)
The ELBO $\hat{\mathcal{L}}(\vartheta)$ can be calculated for any value of $\vartheta$ as every other quantity is known. Hence we can maximize $\hat{\mathcal{L}}(\vartheta)$ over $\vartheta$ to obtain $\hat{\vartheta}\in{\arg\max}_{\vartheta}\;\hat{\mathcal{L}}(\vartheta)$. The estimate of the cut-posterior is then
$\hat{p}_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})=q_{\hat{\vartheta}}(\theta\,|\,\eta)\,p(\eta\,|\,D_{1}).$ (10)
By drawing $Z_{i}\overset{\text{iid}}{\sim}p(Z)$, we can have samples $\{\eta_{i},T_{\hat{\vartheta}}(\eta_{i},Z_{i})\}$ from this cut-posterior. We call our method NeVI-Cut (neural variational inference for cut-Bayes). Note that NeVI-Cut is completely agnostic to what upstream data $D_{1}$ and model were used to obtain the samples $\{\eta_{i}\}$ as long as averages using this sample approximate population averages with respect to $p(\eta\,|\,D_{1})$. It also works with any choice of base distribution, any dimension, and support set of the parameters $\theta$ and $\eta$, and any class of normalizing flows. The next section presents some important classes of normalizing flows that we use in practice. The theoretical justification for their use is presented in Section 4.
3.2 Choices of Flows
We first consider the rational-quadratic neural spline flows (RQ-NSF) (Durkan et al.,, 2019) with two widely used flow architectures: an autoregressive architecture referred to as RQ-NSF(AR) (Papamakarios et al.,, 2017), and a coupling architecture referred to as RQ-NSF(C) (Dinh et al.,, 2016; Kingma and Dhariwal,, 2018).
In RQ-NSF(AR), each transformation $f_{k}$ of the function $T$ is parameterized by an autoregressive neural network. Let $Z=(Z_{[1]},...,Z_{[d]})\in\mathbb{R}^{d}$ denote the base variable. In this architecture, each output dimension $Z_{[i]}^{(k)}$ at layer $k$ only depending on the previous components $z_{[{<i}]}^{(k-1)}$ and the conditional input $\eta$. Specifically, the transformation at flow layer $k$ is defined as
$Z_{[i]}^{(k)}=f_{k}\left(Z_{[i]}^{(k-1)};\phi_{k}\left(Z_{[<i]}^{(k-1)},\eta\right)\right),$
where $f_{k}$ is rational quadratic spline transformation (Durkan et al.,, 2019), and $\phi_{k}$ is a masked conditioner neural network that outputs the spline parameters based on previous latent variables and $\eta$. Since each one-dimensional RQS is strictly monotone, invertibility is guaranteed, and the autoregressive composition preserves this property across all dimensions. The autoregressive structure yields a lower triangular Jacobian, making the log-determinant straightforward to compute. To ensure all coordinates are eventually updated, permutation layers (e.g., reversing or rotating the variable order) are inserted between autoregressive transformations.
RQ-NSF(C) is a coupling-based flow that applies monotonic rational quadratic splines transformation to a subset of variables while leaving others unchanged. For each transformation $f_{k}$, the input $Z^{(k-1)}$ is partitioned into two disjoint subsets $Z_{[\alpha]}^{(k-1)}$ and $Z_{[\beta]}^{(k-1)}$. The variables in $Z_{[\alpha]}^{(k-1)}$ are unchanged, the monotone RQS transformations are applied to $Z_{[\beta]}^{(k-1)}$ according to
$Z_{[\beta]}^{{(k)}}=f_{k}\left(Z_{[\beta]}^{(k-1)};\psi_{k}\left(Z_{[\alpha]}^{(k-1)},\eta\right)\right),$
where $\psi_{k}(\cdot)$ is a neural network that outputs spline parameters (e.g., bin widths, heights, and slopes) conditioned on the retained subset and conditional input, and together $Z^{(k)}=\left(Z_{[\alpha]}^{(k-1)},Z_{[\beta]}^{{(k)}}\right)$ is passed to the next layer. The partition $(\alpha,\beta)$ is specified by a binary mask, following the coupling design of RealNVP (Dinh et al.,, 2016), and alternated or permuted across layers as in Neural Spline Flows (Durkan et al.,, 2019) so that all coordinates are eventually updated. Because the variables in $\alpha$ are unchanged, the Jacobian has a block-triangular form with an identity block for $\alpha$ and a diagonal block for $\beta$, making the log-determinant efficient to compute (Dinh et al.,, 2016).
In addition to spline-based flows, we also consider Unconstrained Monotonic Neural Networks (UMNN; Wehenkel and Louppe,, 2019), which construct invertible transformations by integrating positive neural network outputs. We extend this idea to a conditional setting. Similar to the original UMNN, our construction specifies the variational distribution $q(\theta\,|\,\eta)$ induced by the monotone transformation conditioned on input $\eta$, defined as
$\theta=T_{\vartheta}(\eta,Z)=a_{\vartheta_{a}}(\eta)+\int_{0}^{Z}\exp\left(g_{\vartheta_{g}}(\eta,t)\right)\,dt.$ (11)
Here $a(\cdot)$ is a neural network with parameters $\vartheta_{a}$ that shifts the transformation, and $g(\cdot)$ is a neural network with parameters $\vartheta_{g}$ whose output is passed through an exponential function and models the log-gradient of $T_{\vartheta}$ with respect to $Z$, and $\vartheta=(\vartheta_{a},\vartheta_{g})$. This guarantees that the derivative of $T_{\vartheta}$ with respect to $Z$ is strictly positive and hence that the transformation is monotone and invertible.
3.3 Triply-Stochastic Algorithm
The computational challenges in calculating the objective in (9) can be three-pronged. First, a large number of samples $\eta_{i}$ may be available from the upstream analysis. Second, the expectation with respect to $Z$ is typically not available in closed form. Third, the likelihood $p(D_{2}\,|\,\theta,\eta)$ might be slow to compute especially for large datasets. We address these challenges by leveraging three sources of stochasticity.
Mini-batching or stochastic gradient descent is widely used to speed up neural network optimization. The main idea is that terms in the objective function involving sums or averages over many variables can be replaced by averages (or appropriately scaled sums) over smaller subsets of these variables. These subsets are referred to as mini-batches. Only one batch is used at a time, and the mini-batches are cycled throughout the optimization. Stochastic gradient descent is an extreme case with a mini-batch size of one. Each of these three challenges has a scalable solution using such stochastic approximations. At each iteration, we can replace the average over all $N$ samples of $\eta$ with a small minibatch $\mathcal{B}_{\eta}$ of size $N_{\eta}$. Similarly, the expectation with respect to $Z$ is replaced by an average of $N_{Z}$ iid observations $\{Z_{j}\}$ sampled from $p(Z)$. The doubly stochastic approximation of (9) is thus given by
$\hat{\mathcal{L}}_{DS}(\vartheta)=\frac{1}{N_{\eta}}\frac{1}{N_{Z}}\sum_{i\in\mathcal{B}_{\eta}}\sum_{j=1}^{N_{Z}}\Big[$ $\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{j})|+$ (12)
$\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta_{i},Z_{j}),\eta_{i}\right)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{j})\,|\,\eta_{i}\right)\big].$
Finally, if the downstream data $D_{2}$ consists of iid units $Y_{1},\ldots,Y_{n}$, then we have $p(D_{2}\,|\,\theta,\eta)=\prod_{i=1}^{n}p(Y_{i}\,|\,\theta,\eta)$. And $\log p(D_{2}\,|\,\theta,\eta)$ can be approximated as $\frac{n}{N_{D}}\sum_{k\in\mathcal{B}_{Y}}\log p(Y_{k}\,|\,\theta,\eta)$. This gives the triply stochastic approximation
$\hat{\mathcal{L}}_{TS}(\vartheta)=\frac{1}{N_{\eta}}\frac{1}{N_{Z}}\sum_{i\in\mathcal{B}_{\eta}}\sum_{j=1}^{N_{Z}}\Big[$ $\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{j})|+$ (13)
$\frac{n}{N_{D}}\sum_{k\in\mathcal{B}_{Y}}\log p\big(Y_{k}\,|\,T_{\vartheta}(\eta_{i},Z_{j}),\eta_{i}\big)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{j})\,|\,\eta_{i}\right)\Big].$
Algorithm 1 Neural Variational Inference for Cut-Bayes (NeVI-Cut)
Input: Upstream posterior samples $\{\eta_{i}\}_{i=1}^{N}$, downstream data $D_{2}$, functional form of downstream likelihood $p(D_{2}\,|\,\theta,\eta)$ and downstream prior $p(\theta\,|\,\eta)$, base distribution $p(Z)$, maximum number of iterations $S$, patience $S_{\text{patience}}$ (number of steps before early stopping), neural class $T$ (e.g., Rational Quadratic Splines or Unconstrained Monotonic Neural Network), learning rate $\alpha$.
Initialize: Set some initial values of $\vartheta$
for $s=1$ to $S$ do
  Sample base variables $\{Z_{i}\}_{i=1}^{N}\overset{\text{iid}}{\sim}p(Z)$
  Transform $\theta_{i}=T_{\vartheta}(\eta_{i},Z_{i})$ for $i=1,\ldots,N$
  Estimate Monte Carlo objective:
$\hat{\mathcal{L}}_{Z}(\vartheta)=\frac{1}{M}\sum_{i=1}^{N}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{i})|+\log p\!\left(D_{2}\mid\theta_{i},\eta_{i}\right)+\log p\!\left(\theta_{i}\mid\eta_{i}\right)\big].$
  Update parameters $\vartheta\leftarrow\vartheta+\alpha\,\mathrm{Adam}\bigl(\nabla_{\vartheta}\hat{\mathcal{L}}_{Z}(\vartheta)\bigr)$(gradient ascent)
  Early stopping: End if $\hat{\mathcal{L}}_{Z}(\vartheta)$ fails to improve for $S_{\text{patience}}$ steps
end for
Final estimate: $\hat{\vartheta}\leftarrow\vartheta$
Samples: $\widetilde{Z}_{i}\overset{\text{iid}}{\sim}p(Z);\;\widetilde{\theta}_{i}=T_{\hat{\vartheta}}(\eta_{i},\widetilde{Z}_{i})$ for $i=1,\ldots,N$
Output: Trained conditional flow $\hat{q}(\theta\,|\,\eta)\overset{\text{d}}{=}T_{\hat{\vartheta}}(\eta,Z),Z\sim p(Z)$, samples from the cut-posterior $\{\eta_{i},\widetilde{\theta}_{i}\}_{i=1}^{N}$.
In practice, all three sources are not required simultaneously, and the level of stochasticity, i.e., choices of the minibatch sizes $N_{\eta},N_{Z},N_{D}$ can be adjusted depending on computational trade-offs. For example, if the data are not iid or not large, mini-batching of the likelihood is not recommended, and the third source can be omitted. In our experiments, to ensure comparability with multiple imputation, we restrict stochasticity to Monte Carlo integration over $Z$. We also combine the stochastic approximations of $\eta$ and $Z$, using the fact that the integration is over the product density $(\eta,Z)\sim p(\eta\,|\,D_{1})\,p(Z)$. Hence, for each $\eta_{i}$ we can sample $Z_{i}\overset{\text{iid}}{\sim}p(Z)$ and $\{(\eta_{i},Z_{i})\}$ is a sample from $p(\eta\,|\,D_{1})\,p(Z)$. We then use the objective
$\hat{\mathcal{L}}_{Z}(\vartheta)=\frac{1}{N}\sum_{i=1}^{N}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{i})|+\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta_{i},Z_{i}),\eta_{i}\right)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{i})\,|\,\eta_{i}\right)\big].$ (14)
Optimization is performed using the Adam algorithm (Diederik,, 2014). The overall procedure is summarized in Algorithm 1.
4 Theory
In this section, we provide theoretical results on guarantees on universal approximation of the true joint cut-posterior $p_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})$ by the NeVI-Cut estimator $\hat{p}_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})$ defined as the variational solution in (10), when the conditional variational family $q(\theta\,|\,\eta)$ is modeled using conditional normalizing flows. Our results show that the convergence rates of the estimator depend on the richness of the neural network class used in the conditional flows and the complexity of the true cut-posterior.
4.1 Uniform Kullback-Leibler Approximation Rates of Conditional Normalizing Flows
We first provide results on universal approximation rates (in Kullback-Leibler divergence) of conditional normalizing flows in terms of the richness of the neural network architecture and complexity of the class of conditional densities it is approximating. These results are then used to prove the convergence of NeVI-Cut, but are also of independent importance. Much of the existing theory on fixed target approximation qualities of conditional normalizing flows is about their expressive powers (Huang et al.,, 2018; Papamakarios et al.,, 2017; Battaglia and Nicholls,, 2024) showing that the class of (conditional) normalizing flows can approximate any true (conditional) distribution in terms of weak convergence. When used as variational families, one needs to study properties of the variational solution among the conditional flow class. As the solution is obtained by minimizing the KLD, we need to quantify uniform rates of approximation of flows to a target class of densities in the KLD. This, to our knowledge, has not been developed.
We develop a theory for the setting where the conditioning variable $\eta$ is multivariate, and the observation variable $\theta$ is univariate. This is because autoregressive flows typically specify a sequence of univariate conditionals (i.e., where the observation variable is univariate while the conditioning variable is multivariate). So this setting helps to illustrate the main proof ideas and insights from the theory while balancing the technical details. We first define a suitable class of conditional densities.
Definition 1.
Let $K_{\eta}\in\mathbb{R}^{d_{\eta}}$ be some compact interval, and $K_{M}=[-M,M]\times K_{\eta}$. Define a class $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ of positive, continuous probability densities $p(\theta\,|\,\eta)$ supported on $\mathbb{R}\times K_{\eta}$, characterized by the following properties:
  1. (i)
    Spikiness rate (log-Lipschitz in $\theta$): There is a function $L_{S}:\mathbb{N}\to(0,\infty)$ such that,
    $\big|\log p(\theta\,|\,\eta)-\log p(\theta^{\prime}\,|\,\eta)\big|\;\leq\;L_{S}(M)\,|\theta-\theta^{\prime}|,\;\forall(\theta,\eta),(\theta^{\prime},\eta)\in[-M,M]\times K_{\eta},\;M\in\mathbb{N}.$
  2. (ii)
    Conditionals perturbation rate (log-Lipschitz in $\eta$): There is a function $L_{CP}:\mathbb{N}\to(0,\infty)$ such that,
    $\big|\log p(\theta\,|\,\eta)-\log p(\theta\,|\,\eta^{\prime})\big|\;\leq\;L_{CP}(M)\,\|\eta-\eta^{\prime}\|_{1},\;\forall(\theta,\eta),(\theta,\eta^{\prime})\in[-M,M]\times K_{\eta},\;M\in\mathbb{N}.$
  3. (iii)
    Quantile growth rate: For each $p\in\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ and $\eta\in K_{\eta}$, let
    $F_{p}(\theta\,|\,\eta)=\int_{-\infty}^{\theta}p(t\mid\eta)\,dt,\qquad T_{p}(\eta,z)=F_{p}^{-1}\!\big(\Phi(z)\,|\,\eta\big),\quad z\in\mathbb{R},$ (15)
    with $\Phi$ the standard normal CDF. So $F_{p}(\theta\,|\,\eta)$ is the conditional cdf, so $F^{-1}_{p}(u\,|\,\eta)$ for $u\in[0,1]$ is the conditional quantile map (with respect to the default Unif[0,1][0,1] base) and $T_{p}$ is the conditional quantile map transformed with respect to a standard Normal base. Then,
    1. (a)
      Median bound: There exists $M_{0}\geq 0$ such that $\sup_{\eta\in K_{\eta}}T(\eta,0)=M_{0}<\infty$,
    2. (b)
      Derivative bound: $T(\eta,z)$ is strictly increasing and $C^{1}$ (continuously differentiable) in $z$ and there exists constants $A_{d},B_{d}\geq 0$ such that
      $\sup_{(\eta,z)\in K_{\eta}\times\mathbb{R}}|\log\partial_{z}T(\eta,z)|\leq H(z):=A_{d}+B_{d}|z|.$
  4. (iv)
    Tail thinness lower bound: There exists $\alpha>0$, $c>0$ such that $p(\theta\mid\eta)\geq c\,e^{-\alpha\theta^{2}}$ for all $(\theta,\eta)\in\mathbb{R}\times K_{\eta}$.
Assumptions $(i)$ and $(ii)$ specify that the log conditional density is local Lipschitz on compacts, with respect to both $\theta$ and $\eta$. The log-Lipschitz moduli $L_{S}(M)$ and $L_{CP}(M)$ are nonnegative and non-decreasing in $M$. The modulus $L_{S}(M)$ controls how spiky the conditional density is with respect to the observation variable $\theta$. The parameter $L_{CP}(M)$ controls how smoothly or abruptly the conditional law varies with the conditioner $\eta$. In Assumption $(iii)$, the parameters $M_{0}$, $A_{d}$, $B_{d}$ govern the global spread of the distribution via controlling the conditional quantile map. From the median and derivative bound, we immediately have the following bound on the conditional quantile map:
$|T_{p}(\eta,z)|\;\leq\;G(z):=M_{0}+|z|e^{A_{d}+B_{d}|z|}\quad\forall(\eta,z).$ (16)
Thus, we allow the quantile map can grow upto exponentially faster than the base (Gaussian) quantile, enabling it to model densities with a certain degree of thicker tails than Gaussian. As $M$ grows, most of the probability mass remains within [G(M),G(M)][-G(M),G(M)]. Faster growth of $G(M)$ corresponds to heavier tails (relative to the normal base).
Table 1 summarizes the tail (large $M$) orders of the spikiness modulus $L_{S}(M)$, the conditional perturbation modulus $L_{CP}(M)$, and the quantile growth $G(M)$ for common conditional families whose parameters (e.g., location, scale, shape) depend continuously on $\eta$.
Distribution Spikiness $L_{S}(M)$ Conditional Perturbation Rate $L_{CP}(M)$ Quantile growth $G(M)$
Normal
$N(\mu(\eta),\sigma^{2}(\eta))$ $O(M)$ $O(M^{2})$ $O(M)$
Laplace
$\mathrm{Lap}(\mu(\eta),b(\eta))$ $O(1)$ $O(M)$ $O(M^{2})$
Generalized Gaussian
$\mathrm{GGD}(\mu(\eta),\alpha(\eta),\beta(\eta))$ $O(M^{\max\{0,\beta-1\}})$ $O(M^{\beta}\log M)$ $O(M^{2/\beta})$
Symmetrized Log–Normal
$\tfrac{1}{2}(f_{\mathrm{LN}}(y)+f_{\mathrm{LN}}(-y))$ $O\!\left(Me^{O((\log M)^{2})}\right)$ $O\!\left((\log M)^{2}Me^{O((\log M)^{2})}\right)$ $e^{O(M)}$
Cauchy
$C(\mu(\eta),\gamma(\eta))$ $O(1)$ $O(1)$ $O(Me^{M^{2}/2})$
Student’s $t$
$t_{\mu(\eta),s(\eta),\nu(\eta)}$ $O(1)$
$O(\log M)$ if $\nu$ depends on $\eta$,
$O(1)$ otherwise
$O\!\left((Me^{M^{2}/2})^{1/\nu_{\min}}\right)$,
where $\nu_{\min}:=\inf_{\eta\in K_{\eta}}\nu(\eta)>0$
Finite mixtures of the above $\max$ of component rates $\max$ of component rates $\max$ of component growths
Table 1: Tail orders of $L_{S}(M)$, $L_{CP}(M)$, and $G(M)$ for common distribution families. Distribution parameters are $C^{1}$ in $\eta$. Quantile growths are with respect to the Gaussian base.
These examples span a wide range of behavior under the Gaussian-based quantile map $T(\eta,z)=F_{\eta}^{-1}(\Phi(z))$. For a Gaussian conditional law, the base matches the target: $T(\eta,z)=\mu(\eta)+\sigma(\eta)z$ for $|z|\leq M$, so the quantile growth is linear, $G(M)=O(M)$. The spikiness modulus is also linear, $L_{S}(M)=O(M)$ (the score is linear in $\theta$), while the $\eta$–sensitivity can be $O(M^{2})$ because $\log p$ contains a quadratic term in $\theta$. The generalized Gaussian family interpolates smoothly through Gaussian ($\beta=2$), Laplace ($\beta=1$) via the shape parameter $\beta$. The quantiles grow polynomially in the Gaussian base, with order $G(M)=O(M^{2/\beta})$. The log normal and its symmetrized version have tails that decay more slowly than any Gaussian yet faster than any power law. This leads to Gaussian–base quantiles that grow exponentially, summarized by $G(M)=e^{O(M)}$ and is thus an example which sharply meets the exponential quantile bound (16). Thus, families of Gaussian, Laplace, Generalized Gaussian, and log-normal distributions or their finite linear combinations are subsumed in Definition 1.
In contrast, for Cauchy tails the Gaussian base is much thinner: at $|z|=N$ the Gaussian tail probability is $\varepsilon\asymp\phi(M)/M\approx e^{-M^{2}/2}/(M\sqrt{2\pi})$; inverting the Cauchy tail $(\bar{F}(x)\sim\gamma/(\pi x))$ at that $\varepsilon$ yields a quantile of order $\gamma/(\pi\varepsilon)\asymp Me^{M^{2}/2}$, so $G(M)$ is enormous compared with the Gaussian case, even though $L_{S}(M)=O(1)$ and $L_{CP}(M)=O(1)$. Student-$t$ behaves similarly: $\bar{F}(x)\sim cx^{-\nu}$ implies $G(M)\asymp(Me^{M^{2}/2})^{1/\nu_{\min}}$, while $L_{S}(M)=O(1)$ and $L_{CP}(M)=O(\log M)$. For Laplace, exponential tails give $Q_{\eta}(1-\varepsilon)\sim b\log(1/\varepsilon)$, so with $\varepsilon\asymp\phi(M)/M$ we obtain $G(M)=O(M^{2})$, with $L_{S}(M)=O(1)$ and $L_{CP}(M)=O(M)$. This is not surprising, and it is known that a thin base is a poor choice for thick-tailed distributions (Laszkiewicz et al.,, 2022). In such cases, one should use normalizing flows with thicker thicker-tailed base distribution, and similar theoretical results like we derive here can be obtained.
The triplet $(L_{S},L_{CP},G)$ thus separates local sharpness in $\theta$, sensitivity across $\eta$, and tail spread under the Gaussian base, explaining why heavy-tailed laws exhibit very large $G(M)$ even when their local moduli remain small. The rates are derived in Appendix A. We also note that for any conditional density in the class $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$, Assumption 1(i)(i)$(iii)$ leads to local-Lipschitz bounds on the conditional median and the log-derivative of the conditional quantiles. We state these bounds in the following lemma as they feature in the final approximation rates.
Lemma 1.
Let $p\in\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ and $T_{p}$ denote the corresponding conditional quantile function with a Gaussian base as defined in (15). Define
$\mathsf{L}_{a}$ $=\frac{2}{c}\,e^{\alpha M_{0}^{2}}\,e^{L_{CP}(M_{0})\,\mathrm{diam}(K_{\eta})}\,L_{CP}(M_{0})\,$ (17)
$\mathsf{L}_{b}(M)$ $=\;M\;+\;e^{H(M)}\,L_{S}\!\big(G(M)\big)\;+\;L_{CP}\!\big(G(M)\big),$
Then on $K_{\eta}$, the conditional median $T_{p}(\eta,0)$ is $L_{a}$-Lipschitz and on $K_{M}$, the log-derivative $\log\partial_{z}T_{p}$ is $\mathsf{L}_{b}(M)$ Lipschitz with respect to the $\ell_{1}$ distance $\|(\eta,z)-(\eta^{\prime},z^{\prime})\|_{1}:=\|\eta-\eta^{\prime}\|_{1}+|z-z^{\prime}|$.
Finally, in Assumption 1(iv)(iv) we impose the Gaussian-type lower tail bound
$p(\theta\,|\,\eta)\ \geq\ c\,e^{-\alpha\theta^{2}}\qquad\text{for all }(\theta,\eta)\in\mathbb{R}\times K_{\eta},$
to rule out pathologically thin tails (as the Gaussian tail itself is quite thin). In variational inference, the candidate $q(\theta\,|\,\eta)$ is generally required to have mass wherever the target density has mass. If the target density is too thin for very long stretches, it may require a highly irregular variational approximant with a small tail mass and creates a large spatial variation in the interior to match the target normalizing constant. This destroys uniform control of $\log\!\big(p/q\big)$ and of its modulus of continuity, and breaks the entropy and dominated convergence arguments used to bound the KL error. Further, we emphasize that this is a lower bound on tail thinness and accommodates all heavier tails as long as they satisfy Assumption 1(iii)(iii).
Our first assumption is that the target conditional density follows Assumption 1.
Assumption 1 (True distribution class).
The true density $p(\theta,\eta)\in\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ for some functions positive increasing functions $L_{S}(\cdot),L_{CP}(\cdot):\mathbb{N}\to\mathbb{(}0,\infty)$ and positive constants $M_{0},A_{d},B_{d}$.
Assumption 1 is not restrictive. Any locally log-Lipschitz continuous positive conditional density $p(\theta\,|\,\eta)$ supported on $\mathbb{R}\times K_{\eta}$ belongs to many such classes $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),G(\cdot))$. The smallest $L_{S},L_{CP}$ for such a density $p(\theta\,|\,\eta)$ is induced by its own moduli and growth rates
$L_{S}^{\ast}(M)$ $=\sup_{\eta\in K_{\eta}}\sup_{\begin{subarray}{c}\theta\neq\theta^{\prime}\\ |\theta\,|\,,|\theta^{\prime}|\leq M\end{subarray}}\frac{|\log p(\theta\mid\eta)-\log p(\theta^{\prime}\mid\eta)|}{|\theta-\theta^{\prime}|},$
$L_{CP}^{\ast}(M)$ $=\sup_{|\theta\,|\,\leq M}\sup_{\eta\neq\eta^{\prime}}\frac{|\log p(\theta\mid\eta)-\log p(\theta\mid\eta^{\prime})|}{\|\eta-\eta^{\prime}\|_{1}},$
both of which exist because of local Lipschitz continuity and positivity. So the only real restriction is the median and quantile derivative bounds, which essentially implies the quantile growth rate bound to not be super exponentially growing with respect to the base (Normal) distribution. As discussed above, this is well known in the literature on normalizing flows, and one can simply use a thicker-tailed base distribution if the assumption is violated with respect to the Gaussian base.
Instead of studying the properties of our method (and in general of variational conditional flows) for individual densities, this framework serves to study entire groups of conditional families characterized by their spikiness, conditional perturbation rate, and quantile growth. This is much like how Hölder or Sobolev classes provide functional envelopes for smoothness and decay and are used to derive the theory of functional estimation for these classes rather than individual functions.
We first state a simple identity relating conditional density to the conditional quantile map which is used at multiple places in the theory.
Lemma 2.
Let $p(\theta,\eta)$ denote a joint density. Define the conditional quantile map $T_{p}$ as in (15). Then $T_{p}$ is continuous on $K_{\eta}\times\mathbb{R}$, strictly increasing and continuously differentiable in $z$, and
$\log\partial_{z}T_{p}(\eta,z)\;=\;\log\phi(z)\;-\;\log p\!\left(T_{p}(\eta,z)\mid\eta\right).$ (18)
As the KLD features the log-density, Lemma 2 helps write the KLD in terms of the derivative of the log-quantile function and the log base density $\phi(z)$, enabling us to apply respective bounds for each term to bound the KLD.
Our second assumption is on the class of normalizing flows used as the variational family.
Assumption 2 (Neural Network Flow Classes).
For each integer $m\geq 1$, let
$\mathcal{F}_{m}\;:=\;\{\,T_{\vartheta}:K_{\eta}\times\mathbb{R}\to\mathbb{R}\;:\;\vartheta\in\Theta_{m}\,\}$
where $\Theta_{m}\subset\mathbb{R}^{m}$ is compact. These families satisfy:
$(i)$ Regularity and monotonicity. For every $m$ and $\vartheta\in\Theta_{m}$, $T_{\vartheta}$ is continuous in $(\eta,z)$, is $C^{1}$ in $z$, and is strictly increasing in $z$ on compacts.
$(ii)$ Global linear envelopes. There exists positive constants $M^{*},A^{*},B^{*}$such that
$\sup_{m\geq 1}\ \sup_{\vartheta\in\Theta_{m}}\ \sup_{\eta\in K_{\eta}}\ |T_{\vartheta}(\eta,0)|\ \leq\ M^{*},\qquad\sup_{m\geq 1}\ \sup_{\vartheta\in\Theta_{m}}\ \sup_{(\eta,z)\in K_{\eta}\times\mathbb{R}}\ |\log\partial_{z}T_{\vartheta}(\eta,z)|\ \leq\ A^{*}+B^{*}|z|.$
$(iii)$ Universal approximation. For any conditional density $p\in\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$, let $T_{p}$ be as in (15). Then for every $\varepsilon>0$, there exists an integer $m=C(M,\varepsilon)$ and $\vartheta\in\Theta_{m}$ such that for all $p\in\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$
$\sup_{\eta\in K_{\eta}}\big|T_{\vartheta}(\eta,0)-T_{p}(\eta,0)\big|\leq\varepsilon,\quad\sup_{(\eta,z)\in K_{M}}\big|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\big|\leq\varepsilon.$
We will show in Theorems 2 and 3 that popular conditional flow specifications satisfy Assumption 2. Before that, we provide a more general result on KL approximation rates for any class of conditional flows satisfying the assumption.
Theorem 1 (Uniform KL rates of conditional flows).
Let $q_{T}$ be a class of conditional densities corresponding to the conditional distribution $\theta\,|\,\eta=T(\eta,Z),Z\sim N(0,1),T\in\mathcal{F}_{m}$ for some class of conditional flows $\mathcal{F}_{m}$ satisfying Assumption 2 and with $M^{*}\geq M_{0},A^{*}\geq A_{0},B^{*}\geq B_{0}$, and complexity budget $C(M,\varepsilon)$, for a class $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$. Define $\widetilde{G}(M):=M^{*}+|M|e^{A^{*}+B^{*}|M|}$. Then for any
$m\geq C^{*}(\varepsilon):=C\left(R\sqrt{-\log\varepsilon},\frac{\varepsilon}{4\widetilde{G}(R\sqrt{-\log\varepsilon})L_{S}(\widetilde{G}(R\sqrt{-\log\varepsilon}))}\right),$
there exists $T\in\mathcal{F}_{m}$ such that
$\sup_{\eta\in K_{\eta}}\mathrm{KL}\,\!\big(q_{T}(\theta\mid\eta)\,\|\,p(\theta\mid\eta)\big)\ \leq\ \varepsilon,$
for a constant $R>0$ depending only on (and increasing in) $M^{*},A^{*},B^{*},\alpha,c$.
Theorem 1 is a result of independence interest. To our knowledge, it is the first fixed-target uniform KL approximation rates for conditional normalizing flows and can be used to understand KLD between any target conditional posterior and the variational solution based on conditional normalizing flows. As discussed in the Introduction, existing theoretical results on this topic are either just on the expressiveness of conditional normalizing flows, or statistical guarantees assuming data growth.
We now show how Theorem 1 applies to popular conditional flow classes, with explicit quantification of the number of parameters needed to reach a given accuracy $\varepsilon$. The next result is for the class of Unconstrained Monotonic Neural Networks (UMNN, Wehenkel and Louppe,, 2019) where the conditional distribution $\theta\,|\,\eta)$ is specified as in (11) with the functions $a$ and $g$ modeled as ReLU networks.
Theorem 2 (Rate for Conditional Unconstrained Monotonic Neural Networks).
Let $\eta\in\mathbb{R}^{d_{\eta}}$, $\mbox{diam}_{\eta}=\mbox{diam}(K_{\eta})$, $p(\theta\,|\,\eta)$ satisfies Assumption 1 and $\theta\,|\,\eta=T(\eta,z)$ is a UMNN (11) on $K_{\eta}\times\mathbb{R}$ where $a$ and $g$ are clipped ReLU networks with parameter sets $\vartheta_{a}$ and $\vartheta_{g}$. Let $q_{T}(\theta\,|\,\eta)$ denote the conditional distribution given by the law $T(\eta,Z)$, $Z\sim N(0,1)$. For any small $\varepsilon>0$, let $M(\varepsilon)=R\sqrt{-\log\varepsilon}$ where $R$ is as in Theorem 1, and define $\widetilde{L}_{S}(M)=2ML_{S}(M)$. Then there is a UMNN $q_{T}$ with
$|a_{\vartheta}|$ $\;\leq\;C_{a}\Bigl(\tfrac{\widetilde{L}_{S}\left(\widetilde{G}(M(\varepsilon))\right)L_{a}\operatorname{diam}_{\eta}}{\varepsilon}\Bigr)^{d_{\eta}}\log\tfrac{L_{S}\left(\widetilde{G}(M(\varepsilon))\right)}{\varepsilon},$
$|g_{\vartheta}|$ $\;\leq\;C_{g}\Bigl(\tfrac{\widetilde{L}_{S}\left(\widetilde{G}(M(\varepsilon))\right)L_{b}\left(M(\varepsilon)\right)\left(\operatorname{diam}_{\eta}+M(\varepsilon)\right)}{\varepsilon}\Bigr)^{d_{\eta}+1}\log\tfrac{\widetilde{L}_{S}\left(\widetilde{G}(M(\varepsilon))\right)}{\varepsilon},$
such that
$\sup_{\eta\in K_{\eta}}\;\mathrm{KL}\,\!\big(q_{T}(\theta\mid\eta)\,\|\,p(\theta\mid\eta)\big)\ \leq\ \varepsilon.$
Here $C_{a}$, $C_{g}$ are some universal constants depending only on $d_{\eta}$, and $L_{a}$ and $L_{b}(\cdot)$ are the local Lipschitz constants defined in Lemma 1.
The next Theorem gives a similar result for Rational Quadratic Neural Spline Flows (also introduced in Section 3.2.
Theorem 3 (Rate for Conditional Rational Quadratic Neural Spline Flows).
Let $\eta\in\mathbb{R}^{d_{\eta}}$, $\operatorname{diam}_{\eta}=\operatorname{diam}(K_{\eta})$, and suppose $p(\theta\mid\eta)$ satisfies Assumption 1. Let $T(\eta,z):K_{\eta}\times\mathbb{R}\to\mathbb{R}$ be a conditional rational–quadratic neural spline flow (RQ-NSF) on equispaced knots in [M,M][-M,M], whose per-bin heights and slope logits are produced by a ReLU network with $m$ parameters. Write $q_{T}(\theta\mid\eta)$ for the law of $T(\eta,Z)$, $Z\sim\mathcal{N}(0,1)$. For any small $\varepsilon>0$, set $M=M(\varepsilon)=R\sqrt{-\log\varepsilon}$ (with $R$ as in Theorem 1), and define
$\begin{split}L_{\mathrm{edge}}(M)\ &=\ \frac{e^{H(M)}}{M}\,L_{CP}\!\big(G(M)\big),\quad\varepsilon_{d}\ =\ \frac{\varepsilon}{4\,M\,e^{H(M)}},\\ K_{\star}\ &=\ \Bigl\lceil\frac{L_{b}(M)\,M}{\varepsilon_{d}}\Bigr\rceil\ =\ \Bigl\lceil\frac{4\,L_{b}(M)\,M^{2}e^{H(M)}}{\varepsilon}\Bigr\rceil.\end{split}$
Then there exists an RQ-NSF $T$ such that
$\sup_{\eta\in K_{\eta}}\mathrm{KL}\,\!\big(q_{T}(\theta\mid\eta)\,\|\,p(\theta\mid\eta)\big)\ \leq\ \varepsilon,$
and its parameter count satisfies
$m\ \leq\ C_{r}\,(4K_{\star}+2)\,A^{\,d_{\eta}}\log A,\quad A\ :=\ \frac{4\,M\,e^{H(M)}\bigl(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\operatorname{diam}(K_{\eta})}{\varepsilon},$
for a constant $C_{r}$ depending only on $d_{\eta}$.
The expression of the number of parameters in both Theorems 2 and 3 has similar terms, with the number increasing with the Lipschitz bounds $G(\sqrt{-\log\varepsilon})L_{S}(G(\sqrt{-\log\varepsilon}))$ and $L_{a}$ or $L_{b}(\sqrt{-\log\varepsilon})$ raised to the exponent $d_{\eta}$. This shows that the number of parameters needed to reach a given error rate grows exponentially with the dimension of the function. This is not surprising and aligns with the theory of ReLU neural networks (Yarotsky,, 2017).
4.2 Theory of NeVI-Cut
We now establish error rates of the NeVI-Cut estimate approximating a target cut-posterior using the general theory of Section 4.1. We first provide guarantees on existence and KL approximation rates of the NeVI-Cut estimate obtained as the variational solution based on the loss (6), which integrates the conditional KL over the exact cut distribution $p(\eta\,|\,D_{1})$. We hide the conditioning datasets as they remain fixed for all the theory and simply use $p_{\text{cut}}(\theta,\eta)$, $p_{\text{cut}}(\theta\,|\,\eta)$ and $p_{\text{cut}}(\eta)$ to denote the joint, conditional, and marginal cut-posteriors. For any conditional distribution $q(\theta\,|\,\eta)$, define the expected conditional KLD to the conditional cut-posterior as
$R(q)\;=\;\int\mathrm{KL}\,\!\big(q(\cdot\mid\eta)\,\|\,p_{\text{cut}}(\cdot\mid\eta)\big)\,p_{\text{cut}}(\eta)\,d\eta.$ (19)
Note that $R(q)$ is also the KLD of the joint density $q(\theta\,|\,\eta)p_{\text{cut}}(\eta)$ to the joint cut-posterior $p_{\text{cut}}(\theta,\eta)$, i.e.,
$R(q)=\mathrm{KL}\,\!\Big(q(\theta\,|\,\eta)p_{\text{cut}}(\eta)\,\|\,p_{\text{cut}}(\theta,\eta)\Big).$ (20)
Corollary 1 (Integral-based NeVI-Cut).
Let the target posterior be $p_{\text{cut}}(\theta,\eta)=p_{\text{cut}}(\eta)p_{\text{cut}}(\theta\,|\,\eta)$ where $p_{\text{cut}}(\eta)$ is a distribution with support within a compact set $K_{\eta}\in\mathbb{R}^{d_{\eta}}$ and $p_{\text{cut}}(\theta\,|\,\eta)$ is in $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$. Let $\mathcal{F}_{m}$ be a class of the conditional quantile functions that satisfies Assumption 2 for the class $\mathcal{P}(L_{S}(\cdot),L_{CP}(\cdot),M_{0},A_{d},B_{d})$ with $M^{*}\geq M_{0},A^{*}\geq A_{d},B^{*}\geq B_{d}$, and complexity budget $C(M,\varepsilon)$. Let $Q_{m}$ be the class of conditional distributions corresponding to $\mathcal{F}_{m}$, i.e.,
$\mathcal{Q}_{m}:=\big\{\,q_{T}(\cdot\mid\eta):=\operatorname{Law}\big(T(\eta,Z)\big),\ Z\sim\mathcal{N}(0,1)\ :\ T\in\mathcal{F}_{m}\,\big\}.$
Then, for any small $\varepsilon>0$, we have the following conclusions:
  1. (a)
    For every $m\geq 1$ a minimizer $\hat{q}_{m}\in\arg\min_{q\in\mathcal{Q}_{m}}R(q)$ exist.
  2. (b)
    Fix any accuracy small $\varepsilon>0$ and let $m\geq C^{*}(\varepsilon)$ be the budget of Theorem 1. Then $R(\hat{q}_{m})\;=\;\min_{q\in\mathcal{Q}_{m}}R(q)\;\leq\varepsilon$.
The result shows that the NeVI-Cut solution, using the integrated KL loss (8), approximates the true cut-posterior up to any error rate $\varepsilon$ for a suitably large neural network architecture. For specific choices of flow like UMNN or RQ-NSF, the complexities given by Theorems 2 and 3 hold. The result is in the KLD, which is stronger than the total-variation or the weak distributional metric. In practice, of course, optimization cannot be done using the exact integration as in (6). We use the $N$ samples from $p_{\text{cut}}(\eta)$ and optimize (7). The following result provides the analogous bounds for this actual NeVI-Cut solution.
Theorem 4 (Sample-based NeVI-Cut).
Consider the setting of Corollary 1. Let $\{\eta^{(r)}\}_{r=1}^{N}$ be samples from a stationary ergodic Markov chain with invariant distribution $p_{\text{cut}}(\eta)$. For $q\in\mathcal{Q}_{m}$ set
$R_{N}(q)=\frac{1}{N}\sum_{r=1}^{N}\mathrm{KL}\,\!\big(q(\cdot\mid\eta^{(r)})\,\|\,p(\cdot\mid\eta^{(r)})\big)$
to be the loss used in NeVI-Cut. Then, for each $N$, the minimizer $\hat{q}_{N,m}\in\arg\min_{q\in\mathcal{\mathcal{Q}}_{m}}R_{N}(q)$ exists. Moreover, for any small $\varepsilon\geq 0$, we have $\lim_{N\to\infty}R(\hat{q}_{N,m})\leq\varepsilon$.
Theorem 4 provides existence guarantee and error bound in the average KLD for the actual NeVI-Cut solution using the sample-based loss. The result only assumes that the samples are from a stationary ergodic Markov chain which covers the iid case (Monte Carlo samples) but also covers the more realistic case where the samples are MCMC posterior samples from an upstream Bayesian analysis.
5 Simulation Study
In this section, we first present a simulation study that supports the theoretical justification of our proposed NeVI-Cut, showing that it can accurately recover both the conditional distribution of the parameter of interest and its marginal distribution. We then introduce two reproducible examples from the literature that are frequently discussed in the context of cutting feedback. Because existing cutting feedback methods require access to the upstream data, for this set of illustrations, we have access to the upstream data. Hence, for NeVI-Cut, we first fit an MCMC to the upstream data to obtain posterior samples and then use these samples in Algorithm 1. Notably, these MCMC samples can be replaced by posterior draws obtained from any alternative posterior estimation method for the upstream parameters. In our experiments, the multiple imputation-based nested MCMC for cut-Bayes and the full Bayes are both implemented in rstan (Stan Development Team,, 2025), while the variational inference-based methods, including Gaussian VA-Cut (Yu et al.,, 2023) and NeVI-Cut, are implemented in Python 3.10.
5.1 Expressiveness of NeVI-Cut
We first illustrate the representational capacity of NeVI-Cut for multimodal conditional distributions. We take $p(\eta\,|\,D_{1})=\text{Gamma}(2,1)$, and consider the downstream conditional posterior $p(\theta\,|\,\eta,D_{2})$ which follows a mixture of normal distributions such that:
$p(\theta\,|\,\eta,D_{2})=\pi(\eta)N\left(\mu_{1}(\eta),\sigma_{1}^{2}\right)+\left(1-\pi(\eta)\right)N\left(\mu_{2}(\eta),\sigma_{2}^{2}\right),$
where $\pi(\eta)=0.2+0.5\,\sigma\big(4(\eta-2)\big)$, $\mu_{1}(\eta)=4\tanh(\eta-1)$, $\mu_{2}(\eta)=-4\tanh(\eta+1)$, with fixed $\sigma_{1}^{2}=\sigma_{2}^{2}=1.5$.
We model $T_{\vartheta}$ in (10) using two different classes of flows: an Unconstrained Monotonic Neural Network (UMNN; Wehenkel and Louppe,, 2019) and a Rational Quadratic Neural Spine Flow with Autoregressive Structure (RQ-NSF(AR); Durkan et al.,, 2019). We call these NeVI-Cut.UMNN and NeVI-Cut.NSF respectively. In NeVI-Cut.UMNN(11), the neural network $a(\cdot)$ is a 5-layer multilayer perceptron (MLP) with LeakyReLU activation, and $g(\cdot)$ is a 7-layer MLP with LeakyReLU activation. The base distribution is standard normal.
Figure 2 compares the NeVI-Cut estimates of conditional cut-posteriors $p(\theta\,|\,\eta=\eta_{0})$ with their true values for different choices of $\eta_{0}$, as well as and marginal cut-posterior $p_{\text{cut}}(\theta)$.
Figure 2: Comparison of NeVI-Cut estimates using RQ-NSF(AR) and UMNN flows with the true distributions when estimating the conditional cut-posterior $p(\theta\,|\,\eta=\eta_{0})$ (left) and marginal cut-posterior $p(\theta)$ (right).
To assess the expressiveness of NeVI-Cut for a variety of density shapes, we consider four different choices of $\eta_{0}$ that lead to four different shapes of the conditioning cut-posterior (left panel). For $\eta=0.99$, the density is unimodal with a slight notch to one side of the mode. For $\eta=1.39$, there is a weak second mode. The remaining two choices represent clearly bimodal densities, with $\eta=2.09$ having two roughly equal models, and $\eta=3.05$ having unequal ones. The marginal cut-posterior (right panel) is also bimodal with unequal models. The results show that NeVI-Cut with either choice of flows (UMNN or NSF) is highly accurate, correctly capturing all conditional and marginal cut-posteriors. This small experiment illustrates the expressiveness of NeVI-Cut to model densities of varied shapes. However, NeVI-Cut.UMNN can be computationally demanding due to the need to evaluate integrals. Since the two choices of flows produce nearly identical results, we use the faster approach, NeVI-Cut.NSF, in all subsequent simulation studies and real-world applications. Henceforth, we refer to this implementation simply as NeVI-Cut.
5.2 Misspecified Prior
We revisit a commonly used illustrative example (Liu et al.,, 2009; Jacob et al.,, 2017; Yu et al.,, 2023) that demonstrates the challenges of full Bayesian inference when there is model misspecification or biased data. This example constitutes a case of a misspecified prior in the downstream analysis. The upstream dataset consists of a small sample $z=(z_{1},\ldots,z_{n_{1}})^{\top}$, where $z_{i}\overset{\text{iid}}{\sim}N(\varphi,1)$ and $\varphi$ is the parameter of interest. The downstream dataset is a large sample $w=(w_{1},\ldots,w_{n_{2}})^{\top}$, with $w_{i}\overset{\text{iid}}{\sim}N(\varphi+\eta,1)$. This dataset has a mean $\varphi+\eta$, where $\eta$ represents a bias parameter. We use the same priors as in prior research, i.e., $\varphi\sim N(0,\delta_{1}^{-1})$ and $\eta\sim N(0,\delta_{2}^{-1})$, with $\delta_{2}$ set to a large value to reflect high prior confidence in the absence of bias. When the true bias is large, this prior leads to a biased posterior under the full Bayesian model due to the overconfident assumption that $\eta\approx 0$.
This simulation setting illustrates how posterior samples from the small upstream dataset can be leveraged to inform inference on the bias parameter $\eta$ for the large dataset. Following previous studies, we set $n_{1}=100$ and $n_{2}=1000$, with true values $\varphi=0$ and $\eta=1$. The inverse variance parameters are set to $\delta_{1}=1$ and $\delta_{2}=100$. For this setting, the full Bayes posterior and the cut-posterior for $\eta$ both have closed-form expressions (Appendix D). We can thus compare the marginal distribution of $\eta$ across four methods: the Full Bayes posterior, true cut-posterior, NeVI-Cut estimate of the cut-posterior, and estimated cut-posterior using the Gaussian variational approximation of Yu et al., (2023) (Gaussian VA-Cut).
We evaluate each method using 95% weighted interval score, Continuous Ranked Probability Score (CRPS; Gneiting and Raftery,, 2007), mean squared error (MSE), and 95% coverage as comparison metrics. CRPS measures the accuracy of the predictive distribution by comparing it to the observed value. The 95% weighted interval score assesses the accuracy of prediction based on the quantiles of the predicted samples.
Table 2: Predictive performance metrics across methods in simulation under a misspecified prior. Lower values of weighted interval score, CRPS, and MSE indicate better performance.
Method 95% Interval Score CRPS MSE 95% Coverage
Full Bayes 0.387 (0.346, 0.480) 0.481 (0.440, 0.574) 0.275 (0.231, 0.378) 0.00
True Cut 0.016 (0.009, 0.090) 0.077 (0.029, 0.214) 0.017 (0.002, 0.072) 0.84
Gaussian VA-Cut 0.062 (0.003, 0.213) 0.092 (0.024, 0.250) 0.017 (0.001, 0.071) 0.35
NeVI-Cut 0.016 (0.009, 0.092) 0.077 (0.028, 0.213) 0.017 (0.001, 0.071) 0.83
Table 2 summarizes the performance of different methods. Full Bayes performs uniformly worse compared to other methods, showing the impact of the highly misspecified prior. Conversely, the true cut-posterior has the best metrics, demonstrating the utility of cut-posterior under downstream model misspecification. The Gaussian VA-cut is definitely an improvement over full Bayes and has similar accuracy of the posterior point-estimate as the true Cut (both having the same MSE). However, in all the remaining three metrics which capture distributional accuracy, it is much worse than the true cut, implying that the estimated shape of the cut-posterior from this method is inaccurate. This is because the Gaussian VA method uses a mean-field approximation that tends to underestimate posterior variance. This leads to poor uncertainty quantification and distributinal accuracy, reflected in the very low coverage, and higher interval scores and CRPS. All four metrics for NeVI-Cut are nearly indistinguishable from those of the true cut, implying the high degree of accuracy with which it approximates the cut-posterior.
5.3 Using Propensity Scores under Outcome Model Misspecification
In this section, we present an example of the utility of cut-posterior for misspecified downstream outcome model (likelihood). Cut-posteriors are particularly useful in observational studies where randomization is not feasible. In causal inference, for example, the propensity score is first modeled and then incorporated into the outcome model to estimate causal effects. Hence, propensity score estimation can be viewed as an upstream analysis. Thus, one can deploy cutting feedback methods to propagate uncertainty in posterior samples of propensity scores into the downstream outcome model. Following the simulation design of Jacob et al., (2017), we first estimate propensity scores and then model the effect of a binary treatment $X$ on a binary outcome $Z$, adjusting for potential confounders $C\in\mathbb{R}^{n\times p}$.
The propensity score model (upstream) is specified as
$\text{logit}\,p(X_{i}=1\mid C_{i})$ $=\theta_{1,0}+\sum_{j=1}^{p}\theta_{1,j}C_{ij},\quad i=1,\ldots,n,$
where $X_{i}=1$ indicates treatment and $X_{i}=0$ indicates control. The fitted probabilities from this model serve as the estimated propensity scores $e_{i}$.
To adjust for confounding in the outcome model (downstream), individuals are stratified into quantile-based groups according to their estimated propensity scores. Let $g_{k}(e_{i})$ be a binary indicator equal to 1 if subject $i$’s propensity score $e_{i}$ falls in the $k$-th stratum (e.g., quintile), and 0 otherwise. The outcome model is
$\text{logit}\,p(Z_{i}=1\mid X_{i},e_{i})$ $=\theta_{2,0}+\theta_{2,1}X_{i}+\sum_{k=2}^{5}\theta_{2,k}g_{k}(e_{i}),\quad i=1,\ldots,n,$
where $\theta_{2,1}$ is the causal parameter of interest, quantifying the effect of the treatment $X_{i}$ on the outcome $Z_{i}$. Normal priors are assigned to all parameters, with variance $\sigma^{2}=800$ for the intercept and $\sigma^{2}=50$ for slopes.
For data generation, we simulate a realistic setting where the true treatment assignment depends on confounders. The treatment model (propensity score model) follows the same structure as the analysis model, with true coefficients $\theta_{1}^{*}=(0,0.1,0.2,0.3,0.4,0.5,0.6)$.
The true outcome model is given by:
$\text{logit}\,p(Z_{i}=1|X_{i},C_{i})=\;$ $\gamma_{0}^{*}X_{i}+\gamma_{1}^{*}C_{i1}+\gamma_{2}^{*}\exp(C_{i2}-1)+\gamma_{3}^{*}C_{i3}$
$+\gamma_{4}^{*}\exp(C_{i4}-1)+\gamma_{5}^{*}|C_{i5}|+\gamma_{6}^{*}|C_{i6}|,\quad i=1,\ldots,n,$
with $\gamma^{*}=(0,0.6,0.5,0.4,0.3,0.2,0.1)$.
Comparing the analysis and data-generating models, the upstream analysis model of the propensity scores (treatment model) is correctly specified, but the outcome model is misspecified. Since $\gamma_{0}^{*}=0$, the true causal effect is null, and we evaluate whether each method can recover this correctly via their estimate of $\theta_{2.1}$.
We obtain marginal posterior samples for $\theta_{2,1}$ under different approaches and compare NeVI-Cut with the full Bayes posterior and the multiple imputation-based nested MCMC for the cut-posterior (Plummer,, 2015). The same metrics as in Table 2 are used for performance evaluation, and the findings are summarized in Table 3. Because the true effect is zero, a well-specified method should yield point estimates close to zero with valid uncertainty quantification. The full Bayes posterior suffers from the outcome model misspecification, leading to biased estimates and poor coverage. In contrast, NeVI-Cut once again attains performance nearly identical to that of nested MCMC, both in terms of accuracy and uncertainty quantification. We also report the runtime for each method and observe a substantial computational advantage for NeVI-Cut, which is nearly 40 times faster than the nested MCMC.
Table 3: Performance comparison in the propensity score simulation. Lower values of interval score, CRPS, and MSE indicate better performance.
Method 95% Interval Score CRPS MSE Time (seconds) 95% Coverage
Full Bayes 0.032 (0.015, 0.154) 0.130 (0.049, 0.350) 0.049 (0.005, 0.190) 22919 (14864, 38509) 0.745
Nested MCMC 0.021 (0.015, 0.062) 0.089 (0.043, 0.261) 0.025 (0.003, 0.125) 6344 (5783, 7199) 0.964
NeVI-Cut 0.021 (0.015, 0.065) 0.089 (0.043, 0.266) 0.025 (0.003, 0.127) 185 (182, 190) 0.964
6 Applications on Real Data
6.1 HPV Data
We revisit the epidemiological example from Plummer, (2015); Jacob et al., (2017); Yu et al., (2023), which examines the relationship between human papillomavirus (HPV) prevalence and the incidence of cervical cancer (Maucort-Boulch et al.,, 2008). Our goal is to incorporate posterior samples of HPV prevalence, obtained from an external analysis, into an outcome model that captures their association with cervical cancer incidence.
In the upstream analysis, posterior samples of HPV prevalence are derived using prevalence data from 13 countries. Let $z=(z_{1},\ldots,z_{13})^{\top}$, where $z_{i}$ denotes the number of individuals infected with high-risk HPV out of a sample of size $n_{i}$ in country $i$. The data are modeled as $z_{i}\sim\text{Binomial}(n_{i},\gamma_{i})$, where $\gamma_{i}$ is the parameter representing the true HPV prevalence in country $i$. Independent Beta priors are placed on each prevalence parameter, $\gamma_{i}\sim\text{Beta}(1,1)$. By conjugacy, the posterior distribution is
$\gamma_{i}|z_{i}\sim\text{Beta}(1+z_{i},1+n_{i}-z_{i}),\quad i=1,\ldots,13.$
Posterior samples of $\gamma_{i}$ are then passed to the second module, which models the relationship between HPV prevalence and cervical cancer incidence. Let $w=(w_{1},\ldots,w_{13})^{\top}$ denote the number of observed cervical cancer cases. The downstream analysis model specifies
$w_{i}\sim\text{Poisson}\left(T_{i}\exp(\eta_{1}+\eta_{2}\gamma_{i})\right),$
where $T_{i}$ is the number of years at follow-up for country $i$, and $\eta=(\eta_{1},\eta_{2})^{\top}$ are regression parameters of interest. A multivariate normal prior $\eta\sim N\left(0,1000I_{2}\right)$ is used. Because the downstream model may be misspecified (cervical cancer rates depend on other individual factors as well), cutting feedback methods are particularly relevant in this setting.
Our target is the joint and marginal posterior distributions of $\eta=(\eta_{1},\eta_{2})^{\top}$. We compare the full Bayes and five different cut-posterior approaches: multiple-imputation-based nested MCMC, Gaussian VA-cut (Yu et al.,, 2023), OpenBUGS-Cut which obtains cut-posteriors via the cut function in OpenBUGS with adaptive Metropolis (1D adapter), the Tempered-Cut algorithm of Plummer, (2015), and NeVI-Cut.
The left panel of Figure 3 presents the joint posterior distributions of $\eta$ under the full Bayes and cut-Bayes models. The right panel shows pairwise comparisons of the marginal densities for the four cut-Bayes methods against the baseline cut method of nested MCMC. As noted in prior research (such as, Plummer,, 2015; Yu et al.,, 2023), the full Bayes and cut-posteriors differ substantially, underscoring the sensitivity of inference to model misspecification. The figures on the right panel show that all methods generally produce a reasonable approximation to the nested MCMC cut-posterior. This is not surprising as the cut-posterior is clearly unimodal and reasonably Gaussian-shaped. Overall, NeVI-Cut produces results that most closely align with nested MCMC. Gaussian-VA-cut and Tempered Cut tend to overestimate the spike in $\eta_{22}$, while OpenBUGS-Cut underestimates the spike in $\eta_{21}$.
(a) Joint posterior of $(\eta_{1},\eta_{2})$
(b) Marginal posterior of $(\eta_{1},\eta_{2})$
Figure 3: Posterior distributions for the second-module parameters in the HPV data analysis. The left panel shows joint posteriors of $(\eta_{1},\eta_{2})$ under full Bayes and different cut-Bayes methods. The right panel showcases marginal posteriors of $(\eta_{1},\eta_{2})$ using different cut-Bayes methods compared with multiple imputation-based nested MCMC.
6.2 Estimating Cause-Specific Mortality Fractions in Mozambique
We now present a global health application of NeVI-Cut to propagate published uncertainty estimates of classifier confusion matrices into downstream calibration (bias correction) of predictions from these classifiers. In many low- and middle-income countries (LMICs), complete diagnostic autopsies are often not feasible (Nichols et al.,, 2018), particularly for children and neonatal deaths. As a result, verbal autopsies (VA), a WHO-standardized interview of the caregiver, are widely used for population-level mortality surveillance (Bassat et al.,, 2017; Menendez et al.,, 2017). VA responses are passed through pre-trained classifiers, called computer-coded verbal autopsy (CCVA) algorithms, which make individual-level causes of death (COD) predictions. These predicted CODs are then aggregated over the entire dataset to estimate the cause-specific mortality fractions (CSMFs), which are the prevalences, i.e., the fraction of deaths in the population attributable to each cause.
Extensive benchmarking against gold-standard data with more comprehensive COD ascertainment has revealed that CCVA algorithms often misclassify individual-level COD. The bias due to this misclassification propagates to any downstream use of these predictions, like aggregating them to obtain CSMF estimates. To address this, a VA-calibration framework has been developed that leverages limited gold-standard data (with labeled COD) to estimate the confusion matrices of the CCVA classifiers, and then uses them to correct for the misclassification bias in downstream analysis to estimate CSMFs (Datta et al.,, 2020; Fiksel et al.,, 2022, 2023; Gilbert et al.,, 2023; Pramanik et al., 2025b, ). VA-calibration has been used to calibrate VA-only data collected in the Countrywide Mortality Surveillance for Action (COMSA) program in Mozambique (COMSA-Mozambique; Macicame et al.,, 2023) and produce national-level CSMF estimates in neonates (0-27 days) and children (1-59 months). This nationwide study has exemplified the utility of large-scale implementation of VA studies, and using the VA data to publish mortality rates and monitor COD distributions to inform health policy and track progress toward the Sustainable Development Goals.
However, the current VA-calibration implementation either models the confusion matrices and CSMFs jointly (full Bayes), or uses uncertainty-quantified estimates of confusion matrices from the upstream analysis to construct a parametric prior for the downstream CSMF estimation (sequential Bayes). Either method is essentially targeting a joint full Bayes posterior of the confusion matrices and the CSMFs. However, the downstream CSMF model is likely to be misspecified for the confusion matrices, as these confusion matrices are estimated based on limited gold-standard data collected in a clinical setting, and are being used to calibrate VA data collected in a community setting. Additionally, the downstream data is unlabeled (VA records only with no COD information from autopsy or other sources) and fundamentally cannot inform about the confusion matrix of the classifier, which is based on true labels. Hence, it is undesirable to have reverse feedback from the VA data to the estimates of confusion matrices.
Using the framework developed in Pramanik et al., 2025b , Pramanik et al., 2025a has recently produced uncertainty quantified estimates of confusion matrices of three major CCVA classifiers – InSilicoVA, InterVA, and EAVA, for 2 age groups – neonates and children, and 8 countries (Bangladesh, Ethiopia, Kenya, Mali, Mozambique, Sierra Leone, South Africa, and an “other”). These 48 (3 CCVA $\times$ 2 age-groups $\times$ 8 countries) estimates were obtained using the high-fidelity labeled data from the Child Health and Mortality Prevention Surveillance (CHAMPS) program (Blau et al.,, 2019). CHAMPS is a unique resource for evaluating the confusion matrices of CCVA algorithms, as it reports both the VA records and a COD using a minimally invasive tissue sampling (MITS). The uncertainty-quantified estimates, provided as posterior samples of the misclassification matrix, are publicly available on GitHub.
The published posterior samples of CCVA confusion matrices enable calibration of any VA study using NeVI-Cut. This allows propagating the uncertainty from the confusion matrix estimates into the downstream calibration, while preventing reverse feedback and not requiring access to the CHAMPS data. In this Section, we calibrate the COMSA VA data ($D_{2}$; available on GitHub) using these posterior samples and NeVI-Cut. Considering an age group with a set of $C$ causes and a specific CCVA algorithm $a$, let $\Phi=(\phi_{ij})$ denote the $C\times C$ confusion (misclassification) matrix where
$\phi_{ij}=p(\text{CCVA predicts cause $j$}\,|\,\text{true (MITS) cause is $i$}).$
The CHAMPS data $D_{1}$ was used to obtain the posterior samples $p(\Phi\,|\,D_{1})$. Let $n$ denote the sample size of the COMSA data, and $v_{j}$ denote the count of cases predicted to have COD $j$ by algorithm $a$. If $q=(q_{1},\ldots,q_{C})^{\top}$ denote the vector of probabilities $q_{j}=p(\text{CCVA $a$ predicts cause $j$})$, then by law of total probability $q=\Phi^{\top}p$ where $p=(p_{1},\ldots,p_{C})^{\top}$ denote the CSMF vector, i.e., $p_{i}=p(\text{cause of a death is }i)$. The downstream model for the COMSA VA data $D_{2}$ is
$v\sim\text{Multinomial}(n,q)=\text{Multinomial}(n,\Phi^{\top}p),$
and the primary goal is to estimate calibrated CSMFs $p$. We use the Dirichlet(1,1,,1)($1,\ldots$,1) prior on $p$, which is uniform on the simplex.
For NeVI-Cut, following notations in Section 3.1, we consider the class of conditional distributions $q_{\vartheta}(p\,|\,\Phi)$ corresponding to the law
$\quad p\,|\,\Phi\sim T_{\vartheta}(\text{vec}(\Phi_{-C}),Z),\quad Z\sim N(0,I_{C}).$
Here $T_{\vartheta}$ is a composition of neural network-based transformations with parameters $\vartheta$, and is conditioned on the vectorized misclassification matrix $\text{vec}(\Phi_{-C})$. The transformation maps samples from a base Gaussian distribution to the simplex through a sequence of conditional normalizing flows, followed by a final stick-breaking transformation that ensures $p$ lies on the probability simplex. When vectorizing $\Phi$, the last column is removed (emphasized by the notation $\Phi_{-C}$) since each row lies on the simplex and one column is therefore redundant.
Figure 4: Posterior distributions of calibrated cause-specific mortality fractions (CSMFs) in neonates (0-27 days) based on VA-only data collected by the Countrywide Mortality Surveillance for Action (COMSA) program in Mozambique. NeVI-Cut effectively captures complex distributional features, including multimodality and skewness, achieving similar performance to nested MCMC.
We compare with posteriors from the full Bayes VA-calibration, the modular (sequential) VA-calibration of Pramanik et al., 2025b , that corresponds to a parametric approximation of the full Bayes posterior, and the uncalibrated point estimate $\widehat{p}=\widehat{q}=v/n$ which is obtained by forcing $\Phi$ to be identity in the downstream model. Among cut-Bayes methods, we use the nested MCMC for benchmarking accuracy and timing. To assess the necessity of non-parametric variational families specified via neural networks in NeVI-Cut, we also consider variational cut-Bayes with parametric variational distribution. Specifically, we use a Dirichlet variational family that assumes a linear relationship among the concentration parameters. Formally, the variational distribution is specified as $q_{\alpha}(p\,|\,\Phi)=\text{Dirichlet}(p\,|\,\Phi^{\top}\alpha)$ (referred to as Parametric Cut). We compare the densities approximated by the above methods using both visualizations and distance metrics. Posterior CSMF samples are first transformed using the centered log-ratio (CLR), and the quadratic Wasserstein distance is computed in CLR space between each method and the multiple imputation.
Figure 5: Left panel: Quadratic Wasserstein distances in centered log-ratio space between each method and nested MCMC across algorithms; NeVI-Cut is consistently closest (smallest distance). Right panel: Runtime comparison of NeVI-Cut vs nested MCMC; NeVI-Cut is  2.5-fold faster.
Density plots for the neonate population is shown in Figure 4, and results for children aged 1-59 months are shown in the Supplement Figure 6. There are multiple notable observations. We see that the full Bayes posterior is substantially different from the cut-posteriors, indicating sensitivity to possible model misspecification. As discussed above, the feedback from unlabeled data to confusion matrices is also not meaningful, suggesting the use of cut-posterior for propagating uncertainty in VA-calibration. Particularly for the neonate population, the cut-posterior densities of the calibrated CSMFs exhibit complex features, including multimodality and skewness. The uncalibrated estimates deviate noticeably from the calibrated estimates obtained using nested MCMC and NeVI-Cut, with the latter two showing very close agreement. The closeness of the NeVI-Cut estimate to the nested MCMC is quantified in Figure 5, which plots the Wasserstein distance of all the methods from the nested MCMC. Across all the settings, NeVI-Cut has a significantly smaller Wasserstein distance compared to others. In contrast, results from the parametric cut deviate from those produced by nested MCMC and NeVI-Cut, illustrating that the parametric variational distributions are not sufficiently expressive. In addition, NeVI-Cut demonstrates computational efficiency compared with nested MCMC, running approximately 2.5 times faster.
7 Discussion
The concept of cutting feedback during uncertainty propagation has long been practiced under various names and using an assortment of statistical techniques; but in recent years, the methodology for cutting feedback has become increasingly principled and broadly applied. This is because in many applications, feedback is not scientifically meaningful, or the downstream model is subject to various forms of misspecification, jeopardizing the validity of traditional Bayesian inference. Our contribution, NeVI-Cut, is an important addition to the inventory of cutting feedback methods as it does not need any access to the upstream data or model. Our approach directly utilizes posterior samples from the upstream analysis and employs a stochastic optimization algorithm. NeVI-Cut also relies on minimal assumptions. It does not assume any parametric variational class, but leverages neural network-based normalizing flows to enhance the expressiveness of the variational family.
Through both simulation studies and real-world data examples, we demonstrate that NeVI-Cut can effectively capture both conditional and marginal cut-posteriors of various shapes. Its estimation accuracy for the cut-posterior is comparable to that of nested MCMC, while substantially reducing computation time. The method is implemented in a publicly available software, which only requires users to input the upstream samples, and the downstream data and model.
An important future methodological extension would be to use NeVI-Cut for cutting feedback in more complex analysis schemes beyond the two-module (upstream-downstream) analysis. Modular Bayesian inference, a generalization of cutting feedback, has been recently formalized Liu and Goudie, (2025). A natural extension would be to make NeVI-Cut compatible with analysis involving multiple modules. Also, while model misspecification often motivates the use of cutting feedback, a critical question remains unexplored: how to determine when feedback should be cut. Developing principled criteria for this decision represents an important direction for future work.
Theoretically, we provide, to our knowledge, some of the first results on uniform KL approximation rates of conditional normalizing flows, which are at the heart of NeVI-Cut. These results offer much stronger conclusions than existing results on universal expressiveness of flows and directly lead to guarantees about the convergence of variational solutions (like the NeVI-Cut estimate) using conditional flows to the target collection of conditional distributions. Also, the results are in the fixed-data regime, targeting the actual cut-posterior rather than some large sample limit. The theory also explicitly links the complexity of the neural network architecture to the closeness of approximation. These results will be of independent importance as conditional flows are used in many different applications, e.g., hyper-parameter amortized Bayesian inference (Battaglia and Nicholls,, 2024). Future theoretical work will aim to generalize to multivariate distributions, different base families, and other flow classes.
Acknowledgements
We thank Dr. Scott Zeger and Dr. Jamie Perin for useful discussions about cutting feedback. We acknowledge the use of chatgpt.com for helping with the literature review, proof development, algorithmic implementation, and improving the writing in some parts. This work was developed with support from the National Institute of Environmental Health Sciences (NIEHS) (grant R01ES033739); Gates Foundation Grants INV-034842 and INV-070577; Johns Hopkins University Institute for Data Intensive Engineering and Science Seed Funding; and Eunice Kennedy Shriver National Institute of Child Health K99 NIH Pathway to Independence Award (1K99HD114884-01A1).
Appendix A Derivations for Table 1: Spikiness, conditional perturbation rate, and quantile growth
Throughout, $K_{\eta}\subset\mathbb{R}$ is compact and all parameter maps (location, scale, degrees of freedom, etc.) depend continuously on $\eta$ and are thus uniformly (free of $N$) bounded. Write
$Q_{\eta}(u)=F_{\eta}^{-1}(u),\qquad T(\eta,z)=Q_{\eta}(\Phi(z)),\qquad G(M)=\sup_{\eta\in K_{\eta},\ |z|\leq M}|T(\eta,z)|.$
The moduli are
$\begin{split}L_{S}(M)=\sup_{\eta\in K_{\eta}}\ \sup_{\begin{subarray}{c}\theta\neq\theta^{\prime}\\ |\theta\,|\,,|\theta^{\prime}|\leq M\end{subarray}}\frac{|\log p(\theta\mid\eta)-\log p(\theta^{\prime}\mid\eta)|}{|\theta-\theta^{\prime}|},\\ L_{CP}(M)=\sup_{|\theta\,|\,\leq M}\ \sup_{\eta\neq\eta^{\prime}}\frac{|\log p(\theta\mid\eta)-\log p(\theta\mid\eta^{\prime})|}{\|\eta-\eta^{\prime}\|_{1}}.\end{split}$
For $G(M)$ we evaluate upper envelopes of $Q_{\eta}$ at the Gaussian tail levels $\varepsilon_{N}:=\bar{\Phi}(M)=1-\Phi(M)\sim\phi(M)/M$. On $K_{\eta}$ (compact), continuous parameter maps are bounded; for $L_{CP}(M)$ we use that $\log p(\theta\mid\eta)$ is smooth in its parameters on compact sets and apply the chain rule with the (finite) Lipschitz moduli of the parameter maps induced by continuity on $K_{\eta}$. [^2] For standard parametric families, the derivatives of $\log p$ with respect to the parameters are bounded on compacts; combined with bounded variation of the parameter maps on $K_{\eta}$, this gives the displayed $L_{CP}(M)$ orders.
Normal $N(\mu(\eta),\sigma^{2}(\eta))$.
$Q_{\eta}(u)=\mu(\eta)+\sigma(\eta)\Phi^{-1}(u)\ \Rightarrow\ T(\eta,z)=\mu(\eta)+\sigma(\eta)z,\quad G(M)=\sup_{\eta}\,|\,\mu(\eta)|+\sup_{\eta}\sigma(\eta)\,N=O(M).$
$\partial_{\theta}\log p=-(\theta-\mu)/\sigma^{2}\ \Rightarrow\ L_{S}(M)=O\!\Big(\frac{M}{\sigma^{2}_{\min}}\Big)=O(M).$
$\partial_{\mu}\log p=\frac{\theta-\mu}{\sigma^{2}},\qquad\partial_{\sigma}\log p=-\frac{1}{\sigma}+\frac{(\theta-\mu)^{2}}{\sigma^{3}}\ \Rightarrow\ L_{CP}(M)=O(M^{2}).$
Cauchy $C(\mu(\eta),\gamma(\eta))$.
For $Z\sim\mathcal{N}(0,1)$, by Mill’s ratio, the upper Gaussian tail satisfies $\varepsilon_{N}=\Phi(-N)\sim\frac{\phi(M)}{M}.$
$p(\theta\mid\eta)=\frac{1}{\pi\,\gamma(\eta)}\,\frac{1}{1+\big((\theta-\mu(\eta))/\gamma(\eta)\big)^{2}}.$
Upper tail of the Cauchy. Let $\bar{F}_{\eta}(x)=\Pr_{\eta}(\Theta>x)$. With $u=(t-\mu)/\gamma$,
$\bar{F}_{\eta}(x)=\int_{x}^{\infty}p(t\mid\eta)\,dt=\frac{1}{\pi}\int_{(x-\mu)/\gamma}^{\infty}\frac{du}{1+u^{2}}=\frac{1}{\pi}\Big(\frac{\pi}{2}-\arctan\frac{x-\mu}{\gamma}\Big).$
As $x\to\infty$, $\frac{\pi}{2}-\arctan x=x^{-1}+O(x^{-3})$, so
$\bar{F}_{\eta}(x)=\frac{\gamma}{\pi(x-\mu)}\big(1+O(x^{-2})\big)\sim\frac{\gamma(\eta)}{\pi x}.$
Therefore, for $\varepsilon\downarrow 0$,
$Q_{\eta}(1-\varepsilon)\ \text{satisfies}\ \varepsilon=\bar{F}_{\eta}\big(Q_{\eta}(1-\varepsilon)\big)\sim\frac{\gamma(\eta)}{\pi\,Q_{\eta}(1-\varepsilon)}\ \Rightarrow\ Q_{\eta}(1-\varepsilon)\sim\frac{\gamma(\eta)}{\pi\,\varepsilon}.$
So, $Q_{\eta}(\Phi(-z))=O(1/\Phi(z))=z\exp(z^{2}/2)$ as $z\to\infty$, giving the bound
$G(M)=O\!\left(\frac{M}{\phi(M)}\right)=O\!\left(Me^{M^{2}/2}\right).$
The log–density derivatives are
$\partial_{\theta}\log p=-\frac{2(\theta-\mu)}{\gamma^{2}+(\theta-\mu)^{2}},\qquad\partial_{\mu}\log p=\frac{2(\theta-\mu)}{\gamma^{2}+(\theta-\mu)^{2}},\qquad\partial_{\gamma}\log p=-\frac{1}{\gamma}+\frac{2(\theta-\mu)^{2}}{\gamma\big(\gamma^{2}+(\theta-\mu)^{2}\big)}.$
These satisfy the uniform bounds
$\big|\partial_{\theta}\log p\big|\leq\frac{1}{\gamma},\qquad\big|\partial_{\mu}\log p\big|\leq\frac{1}{\gamma},\qquad\big|\partial_{\gamma}\log p\big|\leq\frac{1}{\gamma}.$
Therefore, $L_{S}(M)=O(1),\mbox{ and }L_{CP}(M)=O(1).$
Student-$t_{\nu(\eta)}$ with location $\mu(\eta)$, scale $s(\eta)$. Tail $\bar{F}_{\eta}(x)\sim c_{\nu(\eta)}\,x^{-\nu(\eta)}$ yields
$Q_{\eta}(1-\varepsilon)\sim(c_{\nu(\eta)}/\varepsilon)^{1/\nu(\eta)}\ \Rightarrow\ G(M)=O\!\Big(\Big(\frac{M}{\phi(M)}\Big)^{1/\nu_{\min}}\Big)=O\!\big((Me^{M^{2}/2})^{1/\nu_{\min}}\big),$
with $\nu_{\min}:=\inf_{\eta\in K_{\eta}}\nu(\eta)>0$.
$\partial_{\theta}\log p=-\frac{(\nu+1)(\theta-\mu)}{\nu s^{2}+(\theta-\mu)^{2}}\ \Rightarrow\ L_{S}(M)=O(1).$
Derivatives w.r.t. $(\mu,s)$ are bounded on compacts as above; the $\nu$–term contains $\tfrac{1}{2}\log\!\big(1+(\theta-\mu)^{2}/(\nu s^{2})\big)$, hence on [M,M][-M,M] gives $O(\log M)$:
$L_{CP}(M)=O(\log M).$
Laplace $\mathrm{Lap}(\mu(\eta),b(\eta))$.
Density and cdf:
$p(\theta\mid\eta)=\frac{1}{2b}\exp\!\Big(-\frac{|\theta-\mu|}{b}\Big),\qquad F_{\eta}(\theta)=\begin{cases}\tfrac{1}{2}\exp\!\big(\tfrac{\theta-\mu}{b}\big),&\theta<\mu,\\[2.0pt] 1-\tfrac{1}{2}\exp\!\big(-\tfrac{\theta-\mu}{b}\big),&\theta\geq\mu.\end{cases}$
Upper tail and quantile. For $x\to\infty$,
$\bar{F}_{\eta}(x)=1-F_{\eta}(x)=\tfrac{1}{2}\exp\!\Big(-\frac{x-\mu}{b}\Big).$
Hence, for $\varepsilon\downarrow 0$,
$\varepsilon=\bar{F}_{\eta}\!\big(Q_{\eta}(1-\varepsilon)\big)=\tfrac{1}{2}\exp\!\Big(-\frac{Q_{\eta}(1-\varepsilon)-\mu}{b}\Big)\quad\Rightarrow\quad Q_{\eta}(1-\varepsilon)=\mu+b\log\frac{1}{2\varepsilon}.$
With normal base $Z\sim\mathcal{N}(0,1)$, let $\varepsilon_{N}=\Phi(-N)\sim\phi(M)/M$. Then
$Q_{\eta}\!\big(\Phi(M)\big)=\mu+b\!\left(\log M-\log\phi(M)+\log\tfrac{1}{2}\right)=\mu+\frac{b}{2}M^{2}+b\log M+O(1),$
so the growth envelope on $\{\lvert z\rvert\leq M\}$ is
$G(M)=O\!\Big(\log\frac{M}{\phi(M)}\Big)=O(M^{2}).$
Log–density derivatives. For $\theta\neq\mu$,
$\partial_{\theta}\log p=-\frac{\operatorname{sgn}(\theta-\mu)}{b},\qquad\partial_{\mu}\log p=\frac{\operatorname{sgn}(\theta-\mu)}{b},\qquad\partial_{b}\log p=-\frac{1}{b}+\frac{|\theta-\mu|}{b^{2}}.$
Uniform bounds on the $\theta$–truncated set $\{|\theta-\mu|\leq M\}$:
$\big|\partial_{\theta}\log p\big|\leq\frac{1}{b},\qquad\big|\partial_{\mu}\log p\big|\leq\frac{1}{b},\qquad\big|\partial_{b}\log p\big|\leq\frac{1}{b}+\frac{M}{b^{2}}.$
Therefore, if $0<b_{\min}\leq b(\eta)\leq b_{\max}<\infty$ uniformly in $\eta$,
$L_{S}(M)=O(1),\qquad L_{CP}(M)=O(M).$
Symmetrized mollified log–normal.
The log-normal distribution is supported on $\mathbb{R}_{>0}$, hence we consider a symmetrized log-normal. Also the log-density of log-normal have infinitely large derivative near 0, which makes it ill behaved, to address this, we use a smoothed version by convoluting with a smooth distribution representing a small bump at zero (which can be made arbitrarily small).
Let $\mu(\eta),\sigma(\eta)$ be continuous on compact $K_{\eta}$ with $0<\sigma_{\min}\leq\sigma(\eta)\leq\sigma_{\max}<\infty$. Define the symmetrized log–normal density
$f_{\mathrm{SLN}}(\theta\mid\eta)=\tfrac{1}{2}f_{\mathrm{LN}}(|\theta\,|\,\mid\eta),\qquad f_{\mathrm{LN}}(x\mid\eta)=\frac{1}{x\sigma(\eta)\sqrt{2\pi}}\exp\!\Big(-\frac{(\log x-\mu(\eta))^{2}}{2\sigma^{2}(\eta)}\Big),\quad x>0.$
Let $K_{h}$ be a symmetric $C^{\infty}$ mollifier with compact support [h,h][-h,h], $\int K_{h}=1$, and $K_{h}(u)\geq\kappa_{0}>0$ for $|u|\leq h/2$. Define the smoothed symmetric log–normal
$p_{h}(\theta\mid\eta)=(f_{\mathrm{SLN}}(\cdot\mid\eta)*K_{h})(\theta)=\int_{\mathbb{R}}f_{\mathrm{SLN}}(x\mid\eta)\,K_{h}(\theta-x)\,dx.$
Then $p_{h}$ is positive, $C^{\infty}$, symmetric, and equals $f_{\mathrm{SLN}}$ outside a $2h$-neighborhood of 0.
Uniform lower bound on [M,M][-M,M]. For any $|\theta\,|\,\leq M$,
$p_{h}(\theta\mid\eta)\ \geq\ \kappa_{0}\!\int_{|\theta-x|\leq h/2}f_{\mathrm{SLN}}(x\mid\eta)\,dx\ \geq\ \frac{\kappa_{0}}{2}\!\int_{I_{\theta}}f_{\mathrm{LN}}(x\mid\eta)\,dx,\quad I_{\theta}:=[\max\{0,|\theta\,|\,-h/2\},\,|\theta\,|\,+h/2].$
Since $f_{\mathrm{LN}}(\cdot\mid\eta)$ is decreasing for $x$ beyond its mode, the worst case over $|\theta\,|\,\leq M$ is attained at $|\theta\,|\,=N$. Thus, for all $\eta\in K_{\eta}$,
$\inf_{|\theta\,|\,\leq M}p_{h}(\theta\mid\eta)\ \geq\ \frac{\kappa_{0}}{2}\!\int_{N-h/2}^{M+h/2}f_{\mathrm{LN}}(x\mid\eta)\,dx\ \geq\ \frac{\kappa_{0}h}{4}\,\inf_{\eta\in K_{\eta}}\ \inf_{x\in[N-h/2,\,M+h/2]}f_{\mathrm{LN}}(x\mid\eta).$
For $x\geq 1$ and any $\eta$,
$f_{\mathrm{LN}}(x\mid\eta)=\frac{1}{x\,\sigma(\eta)\sqrt{2\pi}}\exp\!\Big(-\frac{(\log x-\mu(\eta))^{2}}{2\sigma^{2}(\eta)}\Big)\ \geq\ \frac{1}{x\,\sigma_{\max}\sqrt{2\pi}}\exp\!\Big(-\frac{d(x)^{2}}{2\sigma_{\min}^{2}}\Big),$
where $d(x):=\operatorname{dist}(\log x,\,[\mu_{\min},\mu_{\max}])=\max\{0,\ \log x-\mu_{\max},\ \mu_{\min}-\log x\}$. Hence for $M\geq 1$,
$c_{h}(M):=\inf_{\eta\in K_{\eta}}\ \inf_{|\theta\,|\,\leq M}p_{h}(\theta\mid\eta)\ \geq\ \frac{\kappa_{0}h}{4(M+h/2)\,\sigma_{\max}\sqrt{2\pi}}\exp\!\left(-\frac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\right).$
Spikiness and conditional perturbation. As before,
$L_{S}(M)=\sup_{\eta}\sup_{|\theta\,|\,\leq M}\big|\partial_{\theta}\log p_{h}(\theta\mid\eta)\big|\ \leq\ \frac{\|K_{h}^{\prime}\|_{L^{1}}}{c_{h}(M)}=O\!\Big(\frac{1}{c_{h}(M)}\Big),\qquad L_{CP}(M)=O\!\Big(\frac{(\log M)^{2}}{c_{h}(M)}\Big),$
since $\partial_{\eta}p_{h}=(\partial_{\eta}f_{\mathrm{SLN}})*K_{h}$ and $|\partial_{\eta}\log f_{\mathrm{SLN}}|\lesssim(\log M)^{2}$ on [M,M][-M,M]. With the compact bump, these become
$L_{S}(M)$ $=O\!\Big((M+1)\exp\!\Big(\tfrac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\Big)\Big)=O(Me^{O(\log M)^{2}}),$
$L_{CP}(M)$ $=O\!\Big((\log M)^{2}(M+1)\exp\!\Big(\tfrac{(\log(M+h/2)-\mu_{\max})^{2}}{2\sigma_{\min}^{2}}\Big)\Big)=O((\log M)^{2}Ne^{O(\log M)^{2}}).$
Median and derivative envelope. Symmetry gives $M_{0}=0$. Compact support of $K_{h}$ gives two–sided tail comparison for large $x$,
$c_{1}\,f_{\mathrm{SLN}}(x-h\mid\eta)\ \leq\ p_{h}(x\mid\eta)\ \leq\ c_{2}\,f_{\mathrm{SLN}}(x+h\mid\eta),$
with $c_{1},c_{2}>0$ depending only on $K_{h}$. Convolution and Fubini give the exact identity
$\bar{F}_{h,\eta}(x)=\int_{\mathbb{R}}K_{h}(t)\,\bar{F}_{\eta}(x-t)\,dt$
where $\overline{F}_{\eta}$ is the survival function for $f_{SLN}$ and $\overline{F}_{h,\eta}$ is that for $p_{h}$. Since $\bar{F}_{\eta}$ is nonincreasing and $t\in[-h,h]$ on the support of $K_{h}$, it follows that
$\bar{F}_{\eta}(x+h)\ \leq\ \bar{F}_{h,\eta}(x)\ \leq\ \bar{F}_{\eta}(x-h)\qquad\text{for all }x\in\mathbb{R}.$
Write the generalized quantile functions
$Q_{\eta}(u):=\inf\{x:\ F_{\eta}(x)\geq u\},\qquad Q_{h,\eta}(u):=\inf\{x:\ F_{h,\eta}(x)\geq u\},\qquad u\in(0,1).$
Equivalently, if $S_{\eta}:=\bar{F}_{\eta}$ and $S^{-1}_{\eta}(y):=\inf\{x:\ S_{\eta}(x)\leq y\}$, then $Q_{\eta}(u)=S^{-1}_{\eta}(1-u)$.
Fix $u\in(0,1)$ and set $x_{u}:=Q_{h,\eta}(u)$. Continuity of $F_{h,\eta}$ implies $\bar{F}_{h,\eta}(x_{u})=1-u$. The survival bounds at $x=x_{u}$ give
[…]
By monotonicity of $S_{\eta}$ and of its generalized inverse,
$x_{u}-h\ \leq\ S^{-1}_{\eta}(1-u)\ \leq\ x_{u}+h.$
Since $S^{-1}_{\eta}(1-u)=Q_{\eta}(u)$, we conclude
$\big|\,Q_{h,\eta}(u)-Q_{\eta}(u)\,\big|\ \leq\ h\qquad\text{for all }u\in(0,1).$
Setting $u=\Phi(z)$ yields
$\big|\,T_{h}(\eta,z)-T(\eta,z)\,\big|\ \leq\ h\qquad\text{for all }z\in\mathbb{R},$
with a constant $h$ that depends only on the support radius of $K_{h}$ and is independent of $\eta$.
For $x>0$,
$f_{\mathrm{LN}}(x;m,\sigma)=\frac{1}{x\,\sigma\sqrt{2\pi}}\exp\!\Big(-\frac{(\log x-m)^{2}}{2\sigma^{2}}\Big).$
Write $x_{z}:=T(\eta,z)=\exp\!\big(m(\eta)+\sigma(\eta)z+O(1)\big)$ as $|z|\to\infty$. Using the tail comparison at $x=T_{h}(\eta,z)=x_{z}+O(1)$,
$\log p_{h}\bigl(T_{h}(\eta,z)\mid\eta\bigr)=-\log f_{\mathrm{SLN}}\bigl(x_{z}+O(1)\mid\eta\bigr)+O(1).$
Since $f_{\mathrm{SLN}}(x\mid\eta)=\tfrac{1}{2}f_{\mathrm{LN}}(x;m(\eta),\sigma(\eta))$ for $x>0$,
$\log f_{\mathrm{SLN}}(x_{z}+O(1)\mid\eta)=\frac{\big(\log x_{z}-m(\eta)+O(x_{z}^{-1})\big)^{2}}{2\sigma(\eta)^{2}}+\log x_{z}+\log\sigma(\eta)+O(1).$
With $\log x_{z}=m(\eta)+\sigma(\eta)z+O(1)$ this becomes
$\log f_{\mathrm{SLN}}(x_{z}+O(1)\mid\eta)=\frac{z^{2}}{2}+\frac{O(1)}{\sigma(\eta)}\,|z|+\log x_{z}+O(1)=\frac{z^{2}}{2}+\sigma(\eta)\,z+O(1),$
uniformly over $\eta\in K_{\eta}$. Therefore
$\log p_{h}\bigl(T_{h}(\eta,z)\mid\eta\bigr)=\frac{z^{2}}{2}+\sigma(\eta)\,z+O(1).$
Since $\log\phi(z)=-\tfrac{1}{2}z^{2}-\tfrac{1}{2}\log(2\pi)$, we obtain
$\log\partial_{z}T_{h}(\eta,z)=\bigl[-\tfrac{1}{2}z^{2}+O(1)\bigr]\;-\;\bigl[\tfrac{1}{2}z^{2}+\sigma(\eta)\,z+O(1)\bigr]=-\,\sigma(\eta)\,z+O(1),$
uniformly in $\eta\in K_{\eta}$ as $|z|\to\infty$. With $\sigma_{\min}\leq\sigma(\eta)\leq\sigma_{\max}$ on $K_{\eta}$, this yields the two–sided bound
$\sigma_{\max}\,|z|-C\ \leq\ \log\partial_{z}T_{h}(\eta,z)\ \leq\ \sigma_{\max}\,|z|+C$
for some constant $C<\infty$ depending only on $K_{h}$, $K_{\eta}$, and the bounds on $m(\eta)$ and $\sigma(\eta)$. Equivalently,
$\bigl|\log\partial_{z}T_{h}(\eta,z)\bigr|\ \leq\ C_{0}+C_{1}\,|z|\qquad(|z|\ \text{large}),$
with $C_{1}$ proportional to $\sup_{\eta\in K_{\eta}}\sigma(\eta)$. This establishes the envelopes in Assumption 1(iii) and in turn implies $G(M)=O(e^{O}(M))$.
Generalized Gaussian $\mathrm{GGD}(\mu(\eta),\alpha(\eta),\beta(\eta))$.
Let $r:=|\theta-\mu|/\alpha$ and assume $\alpha(\eta)\in[\underline{\alpha},\overline{\alpha}]$, $\beta(\eta)\in[\underline{\beta},\overline{\beta}]$ with $\underline{\alpha}>0$, $2\geq\overline{\beta}$, $\underline{\beta}>0$, and $|\mu(\eta)|\leq M_{\mu}$ uniformly on $K_{\eta}$. The density is
$p(\theta\mid\eta)=\frac{\beta}{2\alpha\,\Gamma(1/\beta)}\exp\!\big(-r^{\beta}\big).$
For $x\to\infty$ we have $t-\mu>0$ on $x,\infty)$, so
$\bar{F}_{\eta}(x)=\int_{x}^{\infty}\frac{\beta}{2\alpha\,\Gamma(1/\beta)}\exp\!\Big(-\big(\tfrac{t-\mu}{\alpha}\big)^{\beta}\Big)\,dt.$
With the change of variables $u=\big(\tfrac{t-\mu}{\alpha}\big)^{\beta}$, so that $t=\mu+\alpha\,u^{1/\beta}$ and $dt=\alpha\,\beta^{-1}u^{1/\beta-1}\,du$, we obtain
$\bar{F}_{\eta}(x)=\frac{1}{2\,\Gamma(1/\beta)}\int_{((x-\mu)/\alpha)^{\beta}}^{\infty}u^{1/\beta-1}e^{-u}\,du=\frac{1}{2\,\Gamma(1/\beta)}\,\Gamma\!\Big(\tfrac{1}{\beta},\,\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}\Big),$
where $\Gamma(s,z)=\int_{z}^{\infty}t^{s-1}e^{-t}\,dt$ is the upper incomplete gamma function. As $z\to\infty$, $\Gamma(s,z)\sim z^{s-1}e^{-z}$ (see §6.5.32 in Abramowitz and Stegun,, 1948). Taking $s=1/\beta$ and $z=\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}$ gives
$\bar{F}_{\eta}(x)\sim\frac{1}{2\,\Gamma(1/\beta)}\Big(\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}\Big)^{1/\beta-1}\exp\!\Big(-\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}\Big)=\frac{1}{2\,\Gamma(1/\beta)}\Big(\tfrac{x-\mu}{\alpha}\Big)^{1-\beta}\exp\!\Big(-\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}\Big).$
Hence, as $x\to\infty$,
$\bar{F}_{\eta}(x)\asymp\exp\!\Big(-\big(\tfrac{x-\mu}{\alpha}\big)^{\beta}\Big),$
with the precise leading factor $\tfrac{1}{2\,\Gamma(1/\beta)}\big(\tfrac{x-\mu}{\alpha}\big)^{1-\beta}$.
So the $(1-\varepsilon)$ quantile satisfies
$Q_{\eta}(1-\varepsilon)\sim\alpha\,[\log(1/\varepsilon)]^{1/\beta}.$
At Gaussian tail level $\varepsilon_{N}=\Phi(-N)\sim\phi(M)/M$,
$\log\frac{1}{\varepsilon_{N}}=\tfrac{1}{2}M^{2}+\log M+\log\sqrt{2\pi},$
hence
$Q_{\eta}(1-\varepsilon_{N})\sim\alpha\Big(\tfrac{1}{2}M^{2}\Big)^{1/\beta},\qquad G(M)=\sup_{\eta\in K_{\eta}}Q_{\eta}(1-\varepsilon_{N})=O\!\big(M^{2/\underline{\beta}}\big).$
Log density derivatives. We have
$\log p(\theta\mid\eta)=\log\beta-\log(2\alpha)-\log\Gamma(1/\beta)-r^{\beta}.$
Derivative in $\theta$:
$\partial_{\theta}\log p=-\beta\,\mathrm{sgn}(\theta-\mu)\,\frac{|\theta-\mu|^{\beta-1}}{\alpha^{\beta}}.$
Parameter derivatives:
$\partial_{\mu}\log p=\beta\,\mathrm{sgn}(\theta-\mu)\,\frac{|\theta-\mu|^{\beta-1}}{\alpha^{\beta}},$
$\partial_{\alpha}\log p=-\frac{1}{\alpha}+\beta\,\frac{|\theta-\mu|^{\beta}}{\alpha^{\beta+1}},$
$\partial_{\beta}\log p=\frac{1}{\beta}+\frac{\psi(1/\beta)}{\beta^{2}}-r^{\beta}\log r.$
Definition of $L_{S}(M)$. For $|\theta\,|\,\leq M$ and $\eta\in K_{\eta}$,
$L_{S}(M):=\sup_{\eta,|\theta\,|\,\leq M}|\partial_{\theta}\log p(\theta\mid\eta)|\ \leq\ \overline{\beta}\,\underline{\alpha}^{-\overline{\beta}}\,(M+M_{\mu})^{\overline{\beta}-1}.$
Definition of $L_{CP}(M)$. If $\mu,\alpha,\beta$ are Lipschitz in $\eta$ with constants $L_{\mu},L_{\alpha},L_{\beta}$, then
$L_{CP}(M):=\sup_{\eta,|\theta\,|\,\leq M}\Big(|\partial_{\mu}\log p|\,L_{\mu}+|\partial_{\alpha}\log p|\,L_{\alpha}+|\partial_{\beta}\log p|\,L_{\beta}\Big).$
Using the above bounds,
$L_{CP}(M)=O\!\Big(L_{\mu}M^{\overline{\beta}-1}+L_{\alpha}M^{\overline{\beta}}+L_{\beta}M^{\overline{\beta}}\log M+L_{\alpha}+L_{\beta}\Big).$
Finally, as $\beta(\eta)$ is always less than or equals to two, the tail decay is always as thick as a Gaussian and part(iv) of the assumption is satisfied.
Finite mixtures.
Consider a finite mixture
$p(\theta\mid\eta)=\sum_{k=1}^{K}\pi_{k}(\eta)\,p_{k}(\theta\mid\eta),\qquad\pi_{k}(\eta)\geq 0,\ \sum_{k}\pi_{k}(\eta)=1,$
with component cdfs $F_{k}$, survival functions $\bar{F}_{k}$, and overall survival $\bar{F}=1-F$. For each component we define the envelope directly from the upper tail quantile as
$G_{k}(M)\;=\;\sup_{\eta\in K_{\eta}}\,\big|Q_{k,\eta}(1-\varepsilon_{N})\big|,\qquad\varepsilon_{N}=\Phi(-N)\sim\frac{\phi(M)}{M},$
where $Q_{k,\eta}(u)=F_{k,\eta}^{-1}(u)$ is the component quantile.
At the tail level, the mixture is governed by the component (or set of components) with the heaviest tail. Suppose there is a nonempty index set $H$ such that every component outside $H$ has vanishing relative tail mass compared to some $h\in H$, while components within $H$ have comparable tails. For large $x$, the survival of the mixture is
$\bar{F}(x\mid\eta)=\sum_{k=1}^{K}\pi_{k}(\eta)\,\bar{F}_{k}(x\mid\eta).$
Under the heaviest–tail assumption, the terms with $k\notin H$ vanish relative to those with $k\in H$, and the latter are comparable to each other. Hence
$\bar{F}(x\mid\eta)\ \sim\ \sum_{h\in H}\pi_{h}(\eta)\,\bar{F}_{h}(x\mid\eta)\ \sim\ C(\eta)\,\max_{k}\bar{F}_{k}(x\mid\eta),$
uniformly in $\eta$, where $C(\eta)$ lies between $\min_{h\in H}\pi_{h}(\eta)$ and $\sum_{h\in H}\pi_{h}(\eta)$.
To translate this into quantiles, note that for $\varepsilon>0$, the quantile $Q_{\eta}(1-\varepsilon)$ is defined by
$\bar{F}\!\big(Q_{\eta}(1-\varepsilon)\mid\eta\big)=\varepsilon.$
If $\bar{F}(x\mid\eta)\asymp\max_{k}\bar{F}_{k}(x\mid\eta)$, then the solution $x=Q_{\eta}(1-\varepsilon)$ must be of the same order as the largest among the component quantiles $Q_{k,\eta}(1-\varepsilon)$. Thus, at Gaussian tail levels $\varepsilon_{N}=\Phi(-N)$,
$Q_{\eta}(1-\varepsilon_{N})\ \asymp\ \max_{k}Q_{k,\eta}(1-\varepsilon_{N}),\qquad G(M)\ \asymp\ \max_{k}G_{k}(M).$
Turning to the derivatives, let
$\tau_{k}(\theta,\eta)=\frac{\pi_{k}(\eta)p_{k}(\theta\mid\eta)}{\sum_{j}\pi_{j}(\eta)p_{j}(\theta\mid\eta)},\qquad\sum_{k}\tau_{k}=1.$
Then
$\partial_{\theta}\log p=\sum_{k=1}^{K}\tau_{k}\,\partial_{\theta}\log p_{k},\qquad\partial_{\eta}\log p=\sum_{k=1}^{K}\tau_{k}\big(\partial_{\eta}\log p_{k}+\partial_{\eta}\log\pi_{k}\big).$
It follows that
$L_{S}(M)\ \leq\ \max_{k}M_{\theta,k}(M),\qquad L_{CP}(M)\ \leq\ \max_{k}M_{\eta,k}(M)+\sup_{k}\|\nabla_{\eta}\log\pi_{k}\|_{L^{\infty}(K_{\eta})}.$
Assume the mixing weights $\pi_{k}:K_{\eta}\to(0,1)$ are continuously differentiable and bounded below by $\pi_{\min}>0$. Then
$\nabla_{\eta}\log\pi_{k}(\eta)=\frac{\nabla_{\eta}\pi_{k}(\eta)}{\pi_{k}(\eta)}\quad\Rightarrow\quad\|\nabla_{\eta}\log\pi_{k}\|_{L^{\infty}(K_{\eta})}\leq\frac{\|\nabla_{\eta}\pi_{k}\|_{L^{\infty}(K_{\eta})}}{\pi_{\min}}.$
Consequently, the conditional perturbation modulus for the mixture satisfies
$L_{CP}(M)\ \leq\ \max_{k}M_{\eta,k}(M)\ +\ \frac{1}{\pi_{\min}}\sup_{k}\|\nabla_{\eta}\pi_{k}\|_{L^{\infty}(K_{\eta})}.$
The last term is a constant, giving the result.
Appendix B Proofs
In this Section we provide proofs of the general results on conditional normalizing flows and NeVI-Cut solutions using them. Proofs of the results for specific flow models (Theorems 2 and 3) are in the next Section.
Proof of Lemma 1.
Lipschitz modulus for $\log\partial_{z}T_{p}$. For $(\eta_{i},z_{i})\in K_{M}$ set $\theta_{i}=T_{p}(\eta_{i},z_{i})$. Applying Lemma 2, by the two moduli condition on $K_{\eta}\times[-G(M),G(M)]$ and since $\log\phi$ is $N$ Lipschitz on [M,M][-M,M],
$\big|\log\partial_{z}T_{p}(\eta_{1},z_{1})-\log\partial_{z}T_{p}(\eta_{2},z_{2})\big|$ $\leq\big|\log\phi(z_{1})-\log\phi(z_{2})\big|+\big|\log p(\theta_{1}\mid\eta_{1})-\log p(\theta_{2}\mid\eta_{2})\big|$
$\leq M\,|z_{1}-z_{2}|+L_{S}\!\big(G(M)\big)\,|\theta_{1}-\theta_{2}|+L_{CP}\!\big(G(M)\big)\,\|\eta_{1}-\eta_{2}\|_{1}.$
By the mean value theorem and Definition 1(iii)(b), $|\theta_{1}-\theta_{2}|=\big|\int_{z_{1}}^{z_{2}}\partial_{u}T_{p}(\eta_{1},u)\,du\big|\leq e^{H(M)}|z_{1}-z_{2}|$. Thus
$\big|\log\partial_{z}T_{p}(\eta_{1},z_{1})-\log\partial_{z}T_{p}(\eta_{2},z_{2})\big|\;\leq\;\big(M+e^{H(M)}L_{S}(G(M))\big)|z_{1}-z_{2}|\;+\;L_{CP}(G(M))\,\|\eta_{1}-\eta_{2}\|_{1}.$
Therefore $\log\partial_{z}T_{p}$ is $\mathsf{L}_{b}(M)$ Lipschitz on $K_{M}$ with respect to the $\ell_{1}$ distance $\|(\eta,z)-(\eta^{\prime},z^{\prime})\|_{1}:=\|\eta-\eta^{\prime}\|_{1}+|z-z^{\prime}|$.
Lipschitz modulus for $T_{p}(\eta,0)$ in $\eta$. FOr simplicity of notaiton, we use $F_{\eta}(\cdot)$ to denote the conditional cdf $F_{p}(\cdot\,|\,\eta)$. Fix $\eta_{1},\eta_{2}\in K_{\eta}$ and set $u=\Phi(0)$, $\theta_{i}=T(\eta_{i},0)=F_{\eta_{i}}^{-1}(u)$, and $g=F_{\eta_{2}}^{-1}$. Then
$|\theta_{1}-\theta_{2}|=\big|g(F_{\eta_{2}}(\theta_{1}))-g(u)\big|\leq\Big(\sup_{v\ \text{between}\ F_{\eta_{2}}(\theta_{1})\ \text{and}\ u}|g^{\prime}(v)|\Big)\,\big|F_{\eta_{2}}(\theta_{1})-u\big|.$
[…]
$m_{0}:=\inf_{\eta\in K_{\eta},\,|\theta\,|\,\leq M_{0}}p(\theta\mid\eta)\geq c\,e^{-\alpha M_{0}^{2}}>0,$
hence $|g^{\prime}(v)|\leq 1/m_{0}$. Therefore
$|\theta_{1}-\theta_{2}|\leq m_{0}^{-1}\,|F_{\eta_{2}}(\theta_{1})-F_{\eta_{1}}(\theta_{1})|.$
For $|t|\leq M_{0}$, Definition 1(ii) implies
$|\log p(t\mid\eta_{1})-\log p(t\mid\eta_{2})|\leq L_{CP}(M_{0})\,\|\eta_{1}-\eta_{2}\|_{1},$
hence $p(t\mid\eta_{2})\leq e^{L_{CP}(M_{0})\|\eta_{1}-\eta_{2}\|_{1}}p(t\mid\eta_{1})$ and vice versa. Integrating over $(-\infty,\theta]$ yields
$|F_{\eta_{2}}(\theta)-F_{\eta_{1}}(\theta)|\leq 2\big(e^{L_{CP}(M_{0})\|\eta_{1}-\eta_{2}\|_{1}}-1\big).$
Using $e^{x}-1\leq e^{x_{0}}x$ for $x\in[0,x_{0}]$ with $x_{0}=L_{CP}(M_{0})\,\mathrm{diam}(K_{\eta})$, we obtain
$|F_{\eta_{2}}(\theta)-F_{\eta_{1}}(\theta)|\leq 2e^{L_{CP}(M_{0})\,\mathrm{diam}(K_{\eta})}L_{CP}(M_{0})\,\|\eta_{1}-\eta_{2}\|.$
Combining the last two displays gives
$|T(\eta_{1},0)-T(\eta_{2},0)|\leq\frac{2}{c}\,e^{\alpha M_{0}^{2}}\,e^{L_{CP}(M_{0})\,\mathrm{diam}(K_{\eta})}\,L_{CP}(M_{0})\,\|\eta_{1}-\eta_{2}\|,$
which proves Lipschitz continuity of $T(\cdot,0)$ on $K_{\eta}$.
Proof of Lemma 2.
We have $F_{p}\!\big(T_{p}(\eta,z)\,|\,\eta\big)=\Phi(z).$ Differentiate both sides with respect to $z$ and use the chain rule:
$\partial_{\theta}F_{p}\!\big(T(\eta,z)\,|\,\eta\big)\cdot\partial_{z}T_{p}(\eta,z)=\phi(z).$
Because $\partial_{\theta}F_{p}(\theta\,|\,\eta)=p(\theta\mid\eta)$, substituting $\theta=T_{p}(\eta,z)$ yields
$p\!\big(T(\eta,z)\mid\eta\big)\,\partial_{z}T(\eta,z)=\phi(z).$
Both factors are positive, hence taking logarithms gives the result. ∎
Proof of Theorem 1.
We use the metric $d_{M}$ on conditional laws induced by their Gaussian quantile maps:
$d_{M}(p,q):=\sup_{(\eta,z)\in K_{M}}\max\Big(\,|T_{p}(\eta,0)-T_{q}(\eta,0)|,\;|\log\partial_{z}T_{p}(\eta,z)-\log\partial_{z}T_{q}(\eta,z)|\,\Big).$
Step 1: Core uniform approximation on $K_{M}$. By Assumption 2(iii), for budget $m=C(M,\varepsilon)$ there exists $T\in\mathcal{F}_{m}$ such that
$d_{M}\big(q_{T},p\big)\ =\ \sup_{(\eta,z)\in K_{M}}\max\!\Big(\,|T(\eta,0)-T_{p}(\eta,0)|,\;|\log\partial_{z}T(\eta,z)-\log\partial_{z}T_{p}(\eta,z)|\,\Big)\ \leq\ \varepsilon.$ (21)
Write $\Delta_{0}(\eta):=T(\eta,0)-T_{p}(\eta,0),\qquad\Delta_{\ell}(\eta,z):=\log\partial_{z}T(\eta,z)-\log\partial_{z}T_{P}(\eta,z).$ Then $|\Delta_{0}|\leq\varepsilon$ and $|\Delta_{\ell}|\leq\varepsilon$ on $K_{M}$.
Step 2: Bounding the KL on the core $|z|\leq M$. For each $\eta$,
$\mathrm{KL}\!\big(q_{T}(\cdot\mid\eta)\,\|\,p(\cdot\mid\eta)\big)=\mathbb{E}\Big[\log\partial_{z}T_{p}(\eta,Z)-\log\partial_{z}T(\eta,Z)+\log p\,\!\big(T_{p}(\eta,Z)\mid\eta\big)-\log p\,\!\big(T(\eta,Z)\mid\eta\big)\Big].$
On $|Z|\leq M$, the first difference is $\leq\varepsilon$ by (21). For the second, use the spikiness modulus on the bounded $\theta$–range: on $|z|\leq M$,
$|T(\eta,z)-T_{p}(\eta,z)|\leq|\Delta_{0}(\eta)|+\int_{0}^{|z|}\!\big|e^{\Delta_{\ell}(\eta,u)}-1\big|\,\partial_{u}T_{p}(\eta,u)\,du\leq\varepsilon+\varepsilon\,|z|\,e^{H(M)}\leq\varepsilon\big(1+2M\,e^{H(M)}\big),$
because $|\log\partial_{u}T_{p}|\leq H(M)$ and $|e^{\Delta_{\ell}}-1|\leq e^{\varepsilon}-1\leq(e-1)\varepsilon\leq 2\varepsilon$ for $\varepsilon\leq 1/2$. Note that on $K_{M}$, $|T_{p}(\eta,Z)|\leq G(M):=M_{0}+|M|e^{A_{d}+B_{d}|M|}$ by (16). Similarly, $|T(\eta,Z)|\leq\widetilde{G}(M):=M^{*}+|M|e^{A^{*}+B^{*}|M|}$. As $M^{*}>M_{0}$, $A^{*}>A_{d}$, and $B^{*}>B_{d}+C_{d}$, we have $\widetilde{G}(M)\geq G(M)$ and both $T_{p}(\eta,Z)$ and $T(\eta,Z)$ lie in $\widetilde{G}(M),\widetilde{G}(M)]$. Hence,
$\big|\log p\!\big(T(\eta,z)\mid\eta\big)-\log p\!\big(T_{p}(\eta,z)\mid\eta\big)\big|\leq L_{S}\!\big(\widetilde{G}(M)\big)\,|T(\eta,z)-T_{p}(\eta,z)|\leq\varepsilon\,L_{S}\!\big(\widetilde{G}(M)\big)\,(1+2Me^{H(M)}).$
Integrating $\phi(z)$ over $|z|\leq M$ (mass $\leq 1$) gives the core contribution
$\leq\varepsilon\Big\{1+L_{S}\!\big(\widetilde{G}(M)\big)(1+2Me^{H(M)})\Big\}$
$\leq 2\varepsilon\Big\{1+L_{S}\!\big(\widetilde{G}(M)\big)\widetilde{G}(M)\Big\}.$
Step 3: Gaussian tails $|z|>M$. Using the identity above and Assumption 2(ii),
$\log\partial_{z}T(\eta,z)\leq A^{*}+B^{*}|z|,\qquad|T(\eta,z)|\leq|T(\eta,0)|+\!\int_{0}^{|z|}\!\partial_{u}T(\eta,u)\,du\leq M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}.$
This leads to
$T(\eta,z)^{2}$ $\;\leq\;\Big(M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}\Big)^{2}$
 $\;\leq 2M^{*2}+2e^{2A^{*}}z^{2}e^{2B^{*}|z|}.$
By the tail–thinness lower bound in Definition 1(iv), $\log p\!\big(T(\eta,z)\mid\eta\big)\leq\alpha\,T(\eta,z)^{2}-\log c$. With $\log\phi(z)=-\tfrac{z^{2}}{2}-\tfrac{1}{2}\log(2\pi)$, Definition 1(iv) and Assumption 2(ii) give, for $|z|>M$ and all $\eta$,
$\log\phi(z)-\log p\!\big(T(\eta,z)\mid\eta\big)-\log\partial_{z}T(\eta,z)$ $\leq-\frac{z^{2}}{2}-\frac{1}{2}\log(2\pi)-\log c+\alpha\,T(\eta,z)^{2}+A^{*}+B^{*}|z|$
$\leq C_{1}\;+\;C_{2}\,z^{2}\,e^{2B^{*}|z|}\;+\;B^{*}|z|,$
where the explicit constants are
$C_{1}=A^{*}-\tfrac{1}{2}\log(2\pi)-\log c+2\alpha\,M^{*2},\qquad C_{2}=\alpha 2e^{2A^{*}}.\;$
Then the tail bound is
$\int_{|z|>M}\!\phi(z)\,[\cdots]\,dz$ $\leq\frac{1}{\sqrt{2\pi}}\Big\{C_{1}\,I_{0}(M)\;+\;C_{2}\,I_{2}(M)\;+\;B^{*}\,I_{1}(M)\Big\},$
where
$I_{0}(M):=\!\int_{|z|>M}\!e^{-z^{2}/2}dz,\quad I_{1}(M):=\!\int_{|z|>M}\!|z|e^{-z^{2}/2}dz,\quad I_{2}(M):=\!\int_{|z|>M}\!(1+z^{2})e^{-z^{2}/2+2B^{*}|z|}dz.$
By Mills’ ratio,
$I_{0}(M)\leq\frac{2\,\phi(M)}{M},\qquad I_{1}(M)=2\!\int_{M}^{\infty}\!ze^{-z^{2}/2}dz=2\,e^{-M^{2}/2}.$
For $I_{2}$, Using $(t+\gamma)^{2}\leq 2(t^{2}+\gamma^{2})$, we have
$1+(t+\gamma)^{2}\;\leq\;1+2t^{2}+2\gamma^{2}.$
Hence
$I_{2}(M)=2e^{\gamma^{2}/2}\int_{M-\gamma}^{\infty}\!\bigl(1+(t+\gamma)^{2}\bigr)e^{-t^{2}/2}\,dt\;\leq\;2e^{\gamma^{2}/2}M_{\gamma}\!\int_{M-\gamma}^{\infty}\!(1+t^{2})e^{-t^{2}/2}\,dt,$
where $M_{\gamma}:=\max\{1+2\gamma^{2},\,2\}$.
For $x>0$,
$\int_{x}^{\infty}(1+t^{2})e^{-t^{2}/2}\,dt=\int_{x}^{\infty}e^{-t^{2}/2}\,dt+\int_{x}^{\infty}t^{2}e^{-t^{2}/2}\,dt.$
Mills’ ratio gives
$\int_{x}^{\infty}e^{-t^{2}/2}\,dt\leq\frac{1}{x}e^{-x^{2}/2}\qquad(x>0).$
Also,
$\int_{x}^{\infty}t^{2}e^{-t^{2}/2}\,dt=\Big[-t\,e^{-t^{2}/2}\Big]_{x}^{\infty}+\int_{x}^{\infty}e^{-t^{2}/2}\,dt=x\,e^{-x^{2}/2}+\int_{x}^{\infty}e^{-t^{2}/2}\,dt\leq\Big(x+\frac{1}{x}\Big)e^{-x^{2}/2}.$
Hence
$\int_{x}^{\infty}(1+t^{2})e^{-t^{2}/2}\,dt\leq\Big(x+\frac{2}{x}\Big)e^{-x^{2}/2}.$
For $x\geq 1$, $x+2/x\leq 2(1+x)$, so
$\int_{x}^{\infty}(1+t^{2})e^{-t^{2}/2}\,dt\leq 2(1+x)e^{-x^{2}/2}.$
Taking $x=M-\gamma$ with $M\geq\gamma+1$ gives
$I_{2}(M)\;\leq\;4M_{\gamma}\,e^{\gamma^{2}/2}\,(1+M-\gamma)\,e^{-\frac{1}{2}(M-\gamma)^{2}}\;\leq\;C_{\gamma}\,(1+M)\,e^{-\frac{1}{2}(M-\gamma)^{2}},$
with $C_{\gamma}:=4M_{\gamma}e^{\gamma^{2}/2}$.
Setting $\gamma=2B^{*}$ and $M_{\gamma}\leq 1+2\gamma^{2}+2\leq 3+8B^{*2}$, yields
$I_{2}(M)\;\leq\;\widetilde{C}_{B^{*}}\,(1+M)\,e^{-\frac{1}{2}\,(M-2B^{*})^{2}},\mbox{ where }\widetilde{C}_{B^{*}}=4(3+8B^{*2})e^{2B^{*2}}.$
Putting the pieces together,
$\int_{|z|>M}\!\phi(z)\,[\cdots]\,dz$ $\leq\ \frac{C_{1}}{\sqrt{2\pi}}\frac{2\,e^{-M^{2}/2}}{M}\;+\;\frac{B^{*}}{\sqrt{2\pi}}2e^{-M^{2}/2}\;+\;\frac{C_{2}}{\sqrt{2\pi}}\,\widetilde{C}_{B^{*}}\,(1+M)\,e^{-\frac{1}{2}(M-2B^{*})^{2}}\,$
$\leq\,R(1+M)\exp\!\Big(-\tfrac{1}{2}(M-2B^{*})^{2}\Big)$
after absorbing all constants in $R$ and using the largest term. For this to be less than $\varepsilon^{\prime}/2$, we write
$t:=M-2B^{*},\qquad 1+M=1+2B^{*}+t\leq(1+2B^{*})e^{t}.$
we have
$R(1+M)e^{-\tfrac{1}{2}t^{2}}\;\leq\;R(1+2B^{*})\,e^{t-\tfrac{1}{2}t^{2}}\;=\;R(1+2B^{*})\,e^{-\tfrac{1}{2}(t-1)^{2}+\tfrac{1}{2}}.$
For this to be less than $\varepsilon^{\prime}/2$
$e^{-\tfrac{1}{2}(t-1)^{2}}\ \leq\ \frac{\varepsilon^{\prime}}{2\,R(1+2B^{*})\,e^{1/2}}\quad\Longleftrightarrow\quad(t-1)^{2}\ \geq\ 2\log\!\frac{2e^{1/2}R(1+2B^{*})}{\varepsilon^{\prime}}.$
[…]
Absorbing $2e^{1/2}(1+2B^{*})$ in $R$ we can choose $M=2B^{*}+1+\sqrt{\,2\log\!\frac{R}{\epsilon}}$ for a suitable constant $R$. So, for any small $\varepsilon^{\prime}$, $M$ can be expressed as $O(\sqrt{-\log\varepsilon^{\prime}})$ where the constant depends only on $M^{*},A^{*},B^{*},\alpha$ and $c$. ∎
Proof of Corollary 1.
(a) Existence. Recall that the parameter set $\Theta_{\,m}\subset\mathbb{R}^{m}$ is compact by Assumption 2. For each fixed $(\eta,z)$, the maps $\vartheta\mapsto T_{\vartheta}(\eta,z)$ and $\vartheta\mapsto\partial_{z}T_{\vartheta}(\eta,z)$ are continuous by (Assumption 2(i)), hence the integrand
$H_{\vartheta}(\eta,z):=\log\partial_{z}T_{\vartheta}(\eta,z)-\log p\!\big(T_{\vartheta}(\eta,z)\mid\eta\big)$
is continuous in $\vartheta$. By Assumption 2(ii) and Definition 1(iv),
$H_{\vartheta}(\eta,z)\;\leq\;A^{*}+B^{*}|z|-\log c+\alpha\Big(M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}\Big)^{\!2}\;\leq\;C_{1}+C_{2}(1+z^{2})e^{2B^{*}|z|},$
with $C_{1},C_{2}$ independent of $\vartheta$. Since $\int_{\mathbb{R}}(1+z^{2})e^{2B^{*}|z|}\phi(z)\,dz<\infty$ (complete the square), the envelope is integrable against $\phi(z)p(\eta)$. By dominated convergence, $\vartheta\mapsto R_{N}(q_{T_{\vartheta}})$ is continuous on the compact set $\Theta_{\,m}$, hence it attains its minimum; existence of $\hat{q}_{m}$ follows.
(b) KLD bound. By Theorem 1, given a small $\varepsilon$, for the budget $m:=C^{*}(\varepsilon)$ there exists $T\in\mathcal{F}_{m}$ such that $R_{m}\big(q_{T}\big)\;\leq\;\varepsilon.$ Because $\hat{q}_{m}$ minimizes $R(q)$ over $q\in\mathcal{Q}_{m}$,
$R(\hat{q}_{m})\;\leq\;R(q_{T})\leq\varepsilon.$
This proves the claim. ∎
Proof of Theorem 4.
Since $\mathcal{Q}_{m}=\{q_{T_{\vartheta}}:\vartheta\in\Theta_{m}\}$, with a slight abuse of notation, we can write $R(q)$ as $R(\vartheta)$ and $R_{N}(q)$ as $R_{N}(\vartheta)$ for the corresponding $\vartheta$. As $\Theta_{m}$ is compact and $\vartheta\mapsto R_{N}(q_{\vartheta})$ is continuous, by Weierstrass Extreme Value Theorem, we have $\hat{q}_{N}\in\arg\min_{\mathcal{Q}_{m}}R_{N}(q)$.
Before proceeding further we state and prove a proposition on uniform laws of large numbers for stationary ergodic processes.
Proposition 1 (Uniform ergodic LLN for the empirical KL).
Let $p(\theta\,|\,\eta)$ satisfies Assumption 1 for some compact $K_{\eta}$ and $p(\eta)$ is a distribution supported on $K_{\eta}$. Let $\{\eta^{(r)}\}_{r\geq 1}$ be a stationary ergodic Markov chain with invariant distribution $p(\eta)$. Let $\mathcal{F}_{m}=\{T_{\vartheta}|\;\vartheta\in\Theta_{m}\}$ denote a class satisfying Assumption 2. Let $q_{\vartheta}$ to be the conditional distribution $\theta\,|\,\eta=T_{\vartheta}(\eta,Z),Z\sim N(0,1)$ and
$f_{\vartheta}(\eta)=D_{KL}\!\big(q_{\vartheta}(\theta\mid\eta)\,\|\,p(\theta\mid\eta)\big),\qquad R(\vartheta)=\mathbb{E}[f_{\vartheta}(\eta)],\quad R_{N}(\vartheta)=\frac{1}{N}\sum_{r=1}^{N}f_{\vartheta}(\eta^{(r)}).$
Then $\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|\xrightarrow[N\to\infty]{a.s.}0.$
Proof of Proposition 1.
Envelope. Let $H(\vartheta,\eta,z)=\log\phi(z)-\log p\!\big(T(\eta,z)\mid\eta\big)-\log\partial_{z}T(\eta,z)$ and $f_{\vartheta}(\eta)=\mathbb{E}[H(u,\eta,Z)]$. Following the proof of Corollary 1, $f_{\vartheta}(\eta)\leq\tau$ for some positive $\tau$ free of $\eta$ and for each fixed $\eta$, the map $\vartheta\mapsto f_{\vartheta}(\eta)$ and is continuous on compact $\Theta_{m}$. Define for $\varepsilon>0$
$L_{\varepsilon}(\eta)=\sup\{|f_{\vartheta}(\eta)-f_{\vartheta^{\prime}}(\eta)|:\ \vartheta,\vartheta^{\prime}\in\Theta_{m},\ \|\vartheta-\vartheta^{\prime}\|\leq\varepsilon\}.$
Then $L_{\varepsilon}(\eta)\to 0$ pointwise as $\varepsilon\downarrow 0$, and $0\leq L_{\varepsilon}(\eta)\leq 2\tau$.
For each fixed $\vartheta\in\Theta_{m}$, the map $\eta\mapsto f_{\vartheta}(\eta)$ is measurable (it is an integral of a continuous function dominated by an integrable envelope). Let $D\subset\Theta_{m}$ be a countable dense subset. By continuity of $\vartheta\mapsto f_{\vartheta}(\eta)$ on compact $\Theta_{m}$, for every $\eta$
$L_{\varepsilon}(\eta)=\sup\{\,|f_{\vartheta}(\eta)-f_{\vartheta^{\prime}}(\eta)|:\ \vartheta,\vartheta^{\prime}\in D,\ \|\vartheta-\vartheta^{\prime}\|\leq\varepsilon\,\}.$
The rightmost expression is the supremum of a countable family of measurable functions, hence $L_{\varepsilon}$ is measurable. So $L_{\varepsilon}\in L^{1}(p)$ and by dominated convergence,
$\mathbb{E}\,L_{\varepsilon}(\eta)\longrightarrow 0\qquad(\varepsilon\downarrow 0).$ (22)
Since $\{\eta^{(r)}\}_{r\geq 1}$ is stationary ergodic and $L_{\varepsilon}\in L^{1}(p)$, Birkhoff’s ergodic theorem yields
$\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta^{(r)})\;\xrightarrow[N\to\infty]{a.s.}\;\mathbb{E}_{p(\eta)}[L_{\varepsilon}(\eta)].$
Let $\{\vartheta_{j}\}_{j=1}^{n(\varepsilon)}$ be an $\varepsilon$-net of $\Theta_{m}$. For any $\vartheta\in\Theta_{m}$, choose $j$ with $\|\vartheta-\vartheta_{j}\|\leq\varepsilon$ and write
$\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\big|R_{N}(\vartheta)-R_{N}(\vartheta_{j})\big|+\big|R(\vartheta)-R(\vartheta_{j})\big|.$
$\leq\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta^{(r)})+\mathbb{E}\,L_{\varepsilon}(\eta).$
Taking $\sup_{\vartheta\in\Theta_{m}}$ in the inequality, we get
$\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\max_{1\leq j\leq n(\varepsilon)}\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta^{(r)})+\mathbb{E}\,L_{\varepsilon}(\eta).$
Taking $\limsup_{N\to\infty}$ on both sides and applying the ergodic theorem to the empirical average of $L_{\varepsilon}(\eta^{(r)})$ yields
$\limsup_{N\to\infty}\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\max_{1\leq j\leq n(\varepsilon)}\limsup_{N\to\infty}\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+2\,\mathbb{E}\,L_{\varepsilon}(\eta).$ (23)
For each fixed $j$, the ergodic theorem gives $R_{N}(\vartheta_{j})\to R(\vartheta_{j})$ almost surely, hence the first term on the right side of (23) is zero. Therefore
$\limsup_{N\to\infty}\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|\leq 2\,\mathbb{E}\,L_{\varepsilon}(\eta).$
Letting $\varepsilon\downarrow 0$ and using $\mathbb{E}\,L_{\varepsilon}(\eta)\to 0$ finishes the proof.
Returning to the proof of Theorem 4, by Proposition 1,
$\Delta_{N,m}:=\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(q)-R(q)\big|\xrightarrow[N\to\infty]{a.s.}0\qquad\text{for each fixed }m.$
Let $q_{m}^{\star}\in\arg\min_{\mathcal{Q}_{m}}R(q)$, which exists by Corollary 1. Then, for all $M$,
$R(\hat{q}_{N,m})$ $\leq R_{N}(\hat{q}_{N,m})+\Delta_{N,m}\leq R_{N}(q_{m}^{\star})+\Delta_{N,m}\leq R(q_{m}^{\star})+2\Delta_{N,m}.$
Taking $\limsup_{N\to\infty}$ yields $\limsup_{N\to\infty}R(\hat{q}_{N,m})\leq\inf_{q\in\mathcal{Q}_{m}}R(q)\leq\varepsilon$ by Corollary 1.
Appendix C Proofs for neural flow examples
We now prove the results for specific neural flow classes.
C.1 Common network notation and quantitative bounds
Let $K_{M}=K_{\eta}\times[-M,M]$, $d_{\eta}:=\dim(\operatorname{aff}(K_{\eta}))$, dimension of the affine hull of the set $K_{\eta}$, and $\operatorname{diam}_{\eta}:=\sup_{\eta,\eta^{\prime}\in K_{\eta}}\|\eta-\eta^{\prime}\|_{2}$. All suprema are over $K_{M}$. Write $\|\,\cdot\,\|_{\infty}$ for the max norm. Throughout, clipping uses the one Lipschitz map $\mathrm{clip}(x;a,b):=\min\{b,\max\{a,x\}\}$. Consider a fully connected ReLU network $h_{\vartheta}:\mathbb{R}^{d}\to\mathbb{R}$ with depth (number of layers) $L$ and maximal width $W$. Let $W_{\ell}(\vartheta)$ denote the layer matrices after absorbing biases as an extra column. Set $S(\vartheta):=\max_{\ell}\|W_{\ell}(\vartheta)\|_{\mathrm{op}}$ and $S_{\Pi}(\vartheta):=\prod_{\ell=1}^{L}\|W_{\ell}(\vartheta)\|_{\mathrm{op}}$. Then by chain rule and as derivative of ReLU (where it exists) is less than $1$, we have $\sup_{x}\|\nabla h_{\vartheta}(x)\|\leq S_{\Pi}(\vartheta)\leq S(\vartheta)^{L}$. If $\vartheta,\vartheta^{\prime}$ lie in the same axis aligned box $B,B]^{m}$, then each layer perturbation satisfies $\|W_{\ell}(\vartheta)-W_{\ell}(\vartheta^{\prime})\|_{\mathrm{op}}\leq RW\|\vartheta-\vartheta^{\prime}\|_{\infty}$. A layerwise chain rule with ReLU one Lipschitz gives, for every input ball $\|x\|\leq B$,
$\sup_{\|x\|\leq B}\big|h_{\vartheta}(x)-h_{\vartheta^{\prime}}(x)\big|\;\leq\;L\,W\,B\,\big(RW\big)^{L}\,\|\vartheta-\vartheta^{\prime}\|_{\infty}.$
C.2 Unconstrained Monotonic Neural Networks
Proof of Theorem 2.
It is enough to show that the class satisfy Assumption 2 as the rest follows from Theorem 2. Define
$T_{\vartheta}(\eta,z)$ $=a_{\vartheta}(\eta)+\int_{0}^{z}\exp\!\big(g_{\vartheta}(\eta,t)\big)\,dt,$
$\qquad a_{\vartheta}$ $=\mathrm{clip}(\tilde{a}_{\vartheta},-M^{*},M^{*}),$
$\qquad g_{\vartheta}$ $=\mathrm{clip}(\tilde{g}_{\vartheta},-A^{*}-B^{*}|z|,A^{*}+B^{*}|z|),$
where $\tilde{a}_{\vartheta}$ is a ReLU network on $K_{\eta}$ and $\tilde{g}_{\vartheta}$ is a ReLU network on $K_{M}$. The clips are pointwise in the displayed variables. Since clipping is one Lipschitz, it preserves continuity and does not increase moduli.
For every $(\eta,z)$ one has $\partial_{z}T_{\vartheta}(\eta,z)=\exp\big(g_{\vartheta}(\eta,z)\big)>0$. Continuity in $(\eta,z)$ follows from continuity of $\tilde{a}_{\vartheta}$ and $\tilde{g}_{\vartheta}$ and of the exponential and of the integral in the upper limit. Differentiability in $z$ and continuity of this partial derivative follows directly from the fundamental theorem of calculus, yielding $\partial_{z}T_{\vartheta}(\eta,z)=\exp(g_{\vartheta}(\eta,z))$. Finally, compactness of the parameter set is ensured by clipping the weights to [B,B][-B,B] giving $\Theta_{m}=[-B,B]^{m}$. So, Assumption 2(i) is satisfied.
By definition $|T_{\vartheta}(\eta,0)|=|a_{\vartheta}(\eta)|\leq M^{*}$. Also $|\log\partial_{z}T_{\vartheta}(\eta,z)|=|g_{\vartheta}(\eta,z)|\leq A^{*}+B^{*}|z|$. This is precisely the global linear envelope. So, Assumption 2(ii) is satisfied.
Lastly, we establish part (iii) with an explicit construction. Let the target quantile map $T_{p}$ satisfy Definition 1 on $K_{M}$. Define $u_{1}(\eta):=T_{p}(\eta,0)$ and $u_{2}(\eta,z):=\log\partial_{z}T_{p}(\eta,z)$. The definition yields constants $M_{0}>0$ and $A_{d},B_{d},C_{d}\geq 0$ with
$\sup_{\eta\in K_{\eta}}|u_{1}(\eta)|\leq M_{0},\qquad\sup_{(\eta,z)\in K_{M}}|u_{2}(\eta,z)|\leq A_{d}+B_{d}|z|.$
Choose thresholds $M^{*}>M_{0}$, $A^{*}>A_{d}$, $B^{*}>B_{d}$. Hence if we construct $\tilde{a}_{\vartheta},\tilde{g}_{\vartheta}$ so that
$\sup_{\eta\in K_{\eta}}\big|\tilde{a}_{\vartheta}(\eta)-u_{1}(\eta)\big|\leq\varepsilon,\qquad\sup_{(\eta,z)\in K_{M}}\big|\tilde{g}_{\vartheta}(\eta,z)-u_{2}(\eta,z)\big|\leq\varepsilon,\qquad 0<\varepsilon\leq\varepsilon_{\max},$
with $\varepsilon_{\max}:=\min\{M^{*}-M_{0},\ A^{*}-A_{d},\ B^{*}-B_{d}\}$, then the clips are inactive on $K_{M}$ and $a_{\vartheta}=\tilde{a}_{\vartheta}$, $g_{\vartheta}=\tilde{g}_{\vartheta}$ there. It follows that
$\sup_{\eta\in K_{\eta}}|a_{\vartheta}(\eta)-u_{1}(\eta)|\leq\varepsilon,\qquad\sup_{(\eta,z)\in K_{M}}|g_{\vartheta}(\eta,z)-u_{2}(\eta,z)|\leq\varepsilon.$
To realize $(\mathrm{Apx})$ with quantitative size, note that Definition 1 supplies a Lipschitz moduli $\L _{a}$ for $u_{1}$ on $K_{\eta}$ and $\L _{b}(M)$ for $u_{2}$. We approximate the true maps
$u_{1}(\eta)=T(\eta,0)\quad\text{on }K_{\eta},\qquad u_{2}(\eta,z)=\log\partial_{z}T(\eta,z)\quad\text{on }K_{M}$
with quantitative budgets that depend on $L_{a}$ and $L_{b}(M)$.
Theorem 1 in Yarotsky, (2017) proves that for any $d\geq 1$, a Lipschitz function $f$ on $0,1]^{d}$ with $\|f\|_{\infty}\leq 1$ admits a ReLU network approximation with depth $O(-\log(\varepsilon))$ and number of weights $O\Big(\varepsilon^{-d}[-\log(\varepsilon)]\Big)$ that achieves sup norm error at most $\varepsilon$ on $0,1]^{d}$. As $u_{1}$ is $L_{a}$-Lipschitz on $K_{\eta}$ and $u_{2}$ is $L_{b}(M)$-Lipschitz on $K_{M}$, we map each domain to the unit cube by the affine scaling $x\mapsto(x-x_{0})/\operatorname{diam}$. Lipschitz constants scale by the domain diameter, so the rescaled functions are $L_{a}\,\operatorname{diam}_{\eta}$ and $L_{b}(M)(\operatorname{diam}_{\eta}+M)$ Lipschitz on $0,1]^{d}$ up to absolute constants. Applying the theorem of Yarotsky, (2017) and scaling the error back gives realizations $\tilde{a}_{\vartheta}$ and $\tilde{g}_{\vartheta}$ with
$\|\tilde{a}_{\vartheta}-u_{1}\|_{\infty,K_{\eta}}\leq\varepsilon,\qquad\|\tilde{g}_{\vartheta}-u_{2}\|_{\infty,K_{M}}\leq\varepsilon,$
for depth $O(\log(1/\varepsilon))$, and parameter counts
$\tilde{a}_{\vartheta}\;\leq\;C_{a}\Bigl(\tfrac{L_{a}\,\operatorname{diam}_{\eta}}{\varepsilon}\Bigr)^{d_{\eta}}\log\tfrac{1}{\varepsilon},\qquad\#\tilde{g}_{\vartheta}\;\leq\;C_{g}\Bigl(\tfrac{L_{b}(M)(\operatorname{diam}_{\eta}+M)}{\varepsilon}\Bigr)^{d_{\eta}+1}\log\tfrac{1}{\varepsilon}.$
Here $C_{a}$ and $C_{g}$ are fixed constants depending only on $d_{\eta}$. This completes the verification for UMNN. ∎
C.3 Rational Quadratic Neural Spline Flows
Proof of Theorem 3.
Partition [M,M][-M,M] into $2K$ equal bins of width $w=M/K$ with knots $z_{i}=-M+jw$ for $j=0,\ldots,2K$, so $z_{K}=0$. We will choose $K$ later and maintain $K\geq M$, hence $w\leq 1$. For each $\eta$ the conditioner $\psi_{\vartheta}(\eta)$ outputs raw vectors $b^{(h)}(\eta)\in\mathbb{R}^{2K}$, $v(\eta)\in\mathbb{R}^{2K+1}$, and a scalar offset $c(\eta)$ with $|c(\eta)|\leq B_{0}$. Define heights with a total mass $H_{\mathrm{tot}}>0$,
$h_{i}(\eta)=H_{\mathrm{tot}}\ \mathrm{softmax}\!\big(b^{(h)}(\eta)\big)_{i},\qquad y_{0}:=b(\eta),\quad y_{i+1}:=y_{i}+h_{i}(\eta)\in[b(\eta),\,b(\eta)+H_{\mathrm{tot}}],$
and slopes via clipped logits
$\beta_{i}(\eta)=\mathrm{clip}\!\big(v_{i}(\eta),-A^{*}-B^{*}|z_{i}|,A^{*}+B^{*}|z_{i}|\big),\qquad s_{i}(\eta):=e^{\beta_{i}(\eta)}.$
On a bin $z_{i},z_{i+1}]$ with $\Delta_{i}:=h_{i}/w$ and $t:=(z-z_{i})/w\in[0,1]$, the monotone rational quadratic formula (Durkan et al.,, 2019) gives
$T_{\vartheta}(\eta,z)$ $=y_{i}+h_{i}\,\frac{\Delta_{i}t^{2}+s_{i}t(1-t)}{\Delta_{i}+(s_{i+1}+s_{i}-2\Delta_{i})t(1-t)},$
$\partial_{z}T_{\vartheta}(\eta,z)$ $=\frac{\Delta_{i}^{2}\bigl(s_{i+1}t^{2}+2\Delta_{i}t(1-t)+s_{i}(1-t)^{2}\bigr)}{\bigl(\Delta_{i}+(s_{i+1}+s_{i}-2\Delta_{i})t(1-t)\bigr)^{2}}.$
The numerator of $\partial_{z}T_{\vartheta}(\eta,z)$ is strictly positive for $t\in[0,1]$ since $s_{i},s_{i+1}>0$, so the derivative is positive on each bin. By construction of RQ-NSF, for adjacent bins, the value and the first derivative match at the knots. Extending with linear tails $z\mapsto y_{0}+s_{0}(z-z_{0})$ for $z\leq z_{0}$ and $z\mapsto y_{2K}+s_{2K}(z-z_{2K})$ for $z\geq z_{2K}$ yields a globally $C^{1}$ strictly increasing map. This proves Assumption 2(i).
For $z\in[z_{i},z_{i+1}]$, Appendix A of Durkan et al., (2019) implies
$\min\{s_{i}(\eta),s_{i+1}(\eta)\}\leq\partial_{z}T_{\vartheta}(\eta,z)\leq\max\{s_{i}(\eta),s_{i+1}(\eta)\}.$
The clipping gives $e^{-A^{*}-B^{*}|z_{i}|}\leq s_{i}(\eta)\leq e^{A^{*}+B^{*}|z_{i}|}$. Using $|z_{i}|\leq|z|+w$ and $|z_{i+1}|\leq|z|+w$ with $w\leq 1$, we obtain
$\big|\log\partial_{z}T_{\vartheta}(\eta,z)\big|\leq(A^{*}+B^{*})+B^{*}|z|\qquad\text{for all }(\eta,z).$
Moreover
$|T_{\vartheta}(\eta,0)|=|y_{K}(\eta)|\leq|b(\eta)|+\sum_{i=0}^{2K-1}h_{i}(\eta)\leq B_{0}+H_{\mathrm{tot}}=:M^{*}.$
Choosing $B_{0}\geq M_{0}$ establishes Assumption 2(ii).
Let $T_{p}$ satisfy Definition 1 and set $u(\eta,z):=\log\partial_{z}T_{p}(\eta,z)$. By Lemma 1, on $K_{M}=[-M,M]\times K_{\eta}$ the function $u$ is $L_{b}(M)$–Lipschitz with
$L_{b}(M)=M+e^{H(M)}L_{S}(M)+L_{CP}(M).$
Using the RQS formula with endpoint logits $\beta_{i}(\eta)=u(\eta,z_{i})$ on each bin,
$\sup_{\tau\in[0,w]}\big|\log\partial_{z}T_{\vartheta}(\eta,z_{i}+\tau)-u(\eta,z_{i}+\tau)\big|\leq L_{b}(M)\,w,$
hence
$\sup_{(\eta,z)\in K_{M}}\big|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\big|\leq L_{b}(M)\frac{M}{K}.$
Impose the second inequality in Assumption 2(iii) with tolerance $\varepsilon$ by choosing
$K_{\star}:=\Big\lceil\tfrac{L_{b}(M)\,M}{\varepsilon}\Big\rceil,$
so $K_{\star}\geq M$ and $w\leq 1$.
Let $h_{i}^{\star}(\eta)=\int_{z_{i}}^{z_{i+1}}\partial_{z}T_{p}(\eta,z)\,dz$ and $S(\eta)=\sum_{j}h_{j}^{\star}(\eta)=\int_{-M}^{M}e^{u(\eta,t)}dt$. Define
$b_{i}^{\star}(\eta):=\log h_{i}^{\star}(\eta)-a(\eta),\qquad a(\eta):=\log S(\eta)-\log H_{\rm tot},$
so that $H_{\rm tot}\,\mathrm{softmax}\!\big(b^{\star}(\eta)\big)=h^{\star}(\eta)$.
Moreover, Lemma 1 gives $\mathrm{Lip}_{\eta}(u(\cdot,z))\leq L_{b}(M)$; hence
$|h_{i}^{\star}(\eta)-h_{i}^{\star}(\eta^{\prime})|$ $\leq\int_{z_{i}}^{z_{i+1}}e^{H(M)}\,|u(\eta,t)-u(\eta^{\prime},t)|\,dt\leq w\,e^{H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|,$
$|\log h_{i}^{\star}(\eta)-\log h_{i}^{\star}(\eta^{\prime})|$ $\leq\frac{|h_{i}^{\star}(\eta)-h_{i}^{\star}(\eta^{\prime})|}{\min h_{i}^{\star}}\leq e^{2H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|.$
We have $2Me^{-H(M)}\leq S(\eta)\leq 2Me^{H(M)}$ and $\mathrm{Lip}_{\eta}(S)\leq 2Me^{H(M)}L_{b}(M)$, hence
$|a(\eta)-a(\eta^{\prime})|=|\log S(\eta)-\log S(\eta^{\prime})|\leq e^{2H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|.$
Therefore
$|b_{i}^{\star}(\eta)-b_{i}^{\star}(\eta^{\prime})|\leq 2\,e^{2H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|\quad\Rightarrow\quad\mathrm{Lip}_{\eta}\big(b^{\star}\big)\leq 2\,e^{2H(M)}L_{b}(M).$
Set
$F(\eta)=\big(b^{\star}(\eta),\,v^{\star}(\eta),\,c^{\star}(\eta)\big)\in\mathbb{R}^{4K_{\star}+2},\qquad v_{i}^{\star}(\eta)=u(\eta,z_{i}),\quad c^{\star}(\eta)=T_{p}(\eta,-M).$
So, $b^{\star}$ is $2e^{2H(M)}L_{b}(M)$–Lipschitz and by Lemma 1 $v^{\star}$ is $L_{b}(M)$–Lipschitz. We now establish the Lipschitz constant for $c^{*}$. We start from the quantile identity
$F_{p}\!\big(T_{p}(\eta,z)\mid\eta\big)=\Phi(z),$
which at z=Mz=-M gives $F_{p}(c^{\star}(\eta)\mid\eta)=\Phi(-M)$. Differentiating both sides with respect to $\eta$ and applying the chain rule yields
$\nabla_{\eta}F_{p}(\theta\mid\eta)\big|_{\theta=c^{\star}(\eta)}+p(c^{\star}(\eta)\mid\eta)\,\nabla_{\eta}c^{\star}(\eta)=0,$
hence
$\|\nabla_{\eta}c^{\star}(\eta)\|\leq\frac{\|\nabla_{\eta}F_{p}(\theta\mid\eta)\|_{\theta=c^{\star}(\eta)}}{p(c^{\star}(\eta)\mid\eta)}.$
We bound the numerator $\|\nabla_{\eta}F_{p}(\theta\mid\eta)\|_{\theta=c^{\star}(\eta)}$ using a change of variables along the quantile map. For any $z$,
$p\!\big(T_{p}(\eta,z)\mid\eta\big)\,\partial_{z}T_{p}(\eta,z)=\phi(z),$
so for $\theta=c^{\star}(\eta)=T_{p}(\eta,-M)$ we have
$\nabla_{\eta}F_{p}(\theta\mid\eta)\big|_{\theta=c^{\star}(\eta)}$ $=\int_{-\infty}^{c^{\star}(\eta)}\nabla_{\eta}p(t\mid\eta)\,dt=\int_{-\infty}^{-M}\nabla_{\eta}p\!\big(T_{p}(\eta,z)\mid\eta\big)\,\partial_{z}T_{p}(\eta,z)\,dz$
$=\int_{-\infty}^{-M}p\!\big(T_{p}(\eta,z)\mid\eta\big)\,\nabla_{\eta}\log p\!\big(T_{p}(\eta,z)\mid\eta\big)\,\partial_{z}T_{p}(\eta,z)\,dz$
$=\int_{-\infty}^{-M}\phi(z)\,\nabla_{\eta}\log p\!\big(T_{p}(\eta,z)\mid\eta\big)\,dz.$
By Definition 1(ii), for $|\theta|\leq G(M)$ we have $\|\nabla_{\eta}\log p(\theta\mid\eta)\|\leq L_{CP}(G(M))$. Quantile growth ensures $|T_{p}(\eta,z)|\leq G(|z|)\leq G(M)$ for all $z\in(-\infty,-M]$. Therefore
$\Big\|\nabla_{\eta}F_{p}(\theta\mid\eta)\Big\|_{\theta=c^{\star}(\eta)}\leq\int_{-\infty}^{-M}\phi(z)\,L_{CP}(G(M))\,dz=L_{CP}(G(M))\,\Phi(-M).$
For the denominator, using $\partial_{z}T_{p}(\eta,z)=\phi(z)/p(T_{p}(\eta,z)\mid\eta)$ gives
$p(c^{\star}(\eta)\mid\eta)=\frac{\phi(-M)}{\partial_{z}T_{p}(\eta,-M)}.$
From Assumption 2(ii), $|\log\partial_{z}T_{p}(\eta,z)|\leq H(|z|)$, hence $\partial_{z}T_{p}(\eta,-M)\leq e^{H(M)}$, so
$p(c^{\star}(\eta)\mid\eta)\geq\phi(-M)e^{-H(M)}.$
Combining and applying Mill’s ratio, $c^{\star}$ is $L_{\mathrm{edge}}(M)$–Lipschitz with
$\|\nabla_{\eta}c^{\star}(\eta)\|\leq\frac{L_{CP}(G(M))}{\phi(-M)e^{-H(M)}}=e^{H(M)}\,L_{CP}(G(M))/M=:L_{edge}.$
By the ReLU vector sup–norm approximation theorem (Yarotsky,, 2017), applied to $F=(b^{\star},v^{\star},c^{\star})$ with
$\mathrm{Lip}_{\eta}(F)\ \leq\ 2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\qquad\text{where}\qquad L_{\mathrm{edge}}(M)=\frac{e^{H(M)}}{M}\,L_{CP}\!\big(G(M)\big),$
there exists a conditioner $\psi_{\vartheta}:K_{\eta}\to\mathbb{R}^{4K_{\star}+2}$ such that
$\sup_{\eta\in K_{\eta}}\big\|F(\eta)-\psi_{\vartheta}(\eta)\big\|_{\infty}\ \leq\ \varepsilon_{\mathrm{par}},$
and the number of parameters satisfies
$m\ \leq\ C_{r}\,(4K_{\star}+2)\,A^{d_{\eta}}\log A,\qquad A\ :=\ \frac{\big(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\big)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}.$
From $\|F-\psi_{\vartheta}\|_{\infty}\leq\varepsilon_{\mathrm{par}}$ we get
$\sup_{\eta\in K_{\eta}}|c(\eta)-c^{\star}(\eta)|\ \leq\ \varepsilon_{\mathrm{par}}.$
We now derive choice of $\varepsilon_{\text{par}}$ such that both the log derivative and the median errors are bounded by $\varepsilon$. For the derivative, let
$\delta\coloneqq\sup_{(\eta,z)\in K_{M}}\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|.$
From the RQ interpolation argument, for each bin $z_{i},z_{i+1}]$,
$\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|\leq L_{b}(M)\,w+\tau,\qquad w=\frac{M}{K_{\star}},$
where
$\tau\ :=\ \max_{i}\,\bigl|\beta_{i}(\eta)-v_{i}^{*}(\eta)\bigr|=\max_{i}\bigl|v_{i}(\eta)-v_{i}^{*}(\eta)\bigr|\ \leq\ \varepsilon_{\mathrm{par}}.$
Choose
$K_{\star}=\Bigl\lceil\frac{L_{b}(M)M}{\varepsilon_{d}}\Bigr\rceil\qquad\text{with}\qquad\varepsilon_{d}=\frac{\varepsilon}{4\,M\,e^{H(M)}},$
so that $L_{b}(M)w\leq\varepsilon_{d}$. Then
$\delta\ \leq\ L_{b}(M)w+\tau\ \leq\ \varepsilon_{d}+\varepsilon_{\mathrm{par}}\ \leq\ \frac{\varepsilon}{4Me^{H(M)}}+\frac{\varepsilon}{4Me^{H(M)}}\ =\ \frac{\varepsilon}{2Me^{H(M)}},$
and hence
$\sup_{(\eta,z)\in K_{M}}\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|\ \leq\ \varepsilon.$
For the median bound, note that
$T(\eta,0)=c(\eta)+\int_{-M}^{0}\partial_{z}T(\eta,t)\,dt.$
Therefore
$\sup_{\eta\in K_{\eta}}\bigl|T_{\vartheta}(\eta,0)-T_{p}(\eta,0)\bigr|$ $\leq\|c-c^{\star}\|_{\infty}+\sup_{\eta}\int_{-M}^{0}\bigl|\partial_{z}T_{\vartheta}(\eta,t)-\partial_{z}T_{p}(\eta,t)\bigr|\,dt$
$\leq\varepsilon_{\mathrm{par}}+M\,e^{H(M)}\bigl(e^{\delta}-1\bigr).$
Using $\delta\leq\varepsilon/(2Me^{H(M)})$ and $e^{x}-1\leq x$ for $x\geq 0$ gives $Me^{H(M)}(e^{\delta}-1)\leq\varepsilon/2$, and since $\varepsilon_{\mathrm{par}}\leq\varepsilon/2$,
$\sup_{\eta\in K_{\eta}}\bigl|T_{\vartheta}(\eta,0)-T_{p}(\eta,0)\bigr|\ \leq\ \varepsilon.$
Thus both inequalities in Assumption 2(iii) hold with tolerance $\varepsilon$.
Finally, the parameter budget becomes
$m\ \leq\ C_{r}\bigl(4K_{\star}+2\bigr)\left(\frac{\bigl(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}\right)^{d_{\eta}}$
$\times\log\!\left(\frac{\bigl(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}\right),$
with
$K_{\star}=\Bigl\lceil\frac{L_{b}(M)M}{\varepsilon_{d}}\Bigr\rceil,\qquad\varepsilon_{d}=\frac{\varepsilon}{4\,M\,e^{H(M)}},\qquad\varepsilon_{\mathrm{par}}=\min\!\Big\{\frac{\varepsilon}{2},\ \frac{\varepsilon}{4\,M\,e^{H(M)}}\Big\}.$
Appendix D Full Bayes and exact cut in misspecified prior
Suppose we have the upstream analysis based on a relatively small sample size $n_{1}$, a following downstream analysis with a large sample size $n_{2}$, the model is as follows:
$z_{i}\overset{\text{iid}}{\sim}N(\phi,1),\quad i=1,\dots,n_{1},$
$w_{i}\,|\,\phi,\eta\overset{\text{iid}}{\sim}N(\phi+\eta,1),\quad i=1,\dots,n_{2},$
with Gaussian priors
$\phi\sim N(0,\delta_{1}^{-1}),\qquad\eta\sim N(0,\delta_{2}^{-1}).$
Let $S_{z}=\sum_{i=1}^{n_{1}}z_{i},S_{w}=\sum_{i=1}^{n_{2}}w_{i}$.
Full Bayes posterior
The joint distribution of $(\phi,\eta)$ is given by
$p(\phi,\eta\,|\,z,w)\,$ $\propto\,p(z\,|\,\phi)\,p(w\,|\,\phi,\eta)\,p(\phi)\,p(\eta)\,$
$\propto\,\exp\left(-\frac{1}{2}\left[\sum_{i=1}^{n_{1}}(z_{i}-\phi)^{2}+\sum_{i=1}^{n_{2}}(w_{i}-\phi-\eta)^{2}\right]-\frac{\delta_{1}}{2}\phi^{2}-\frac{\delta_{2}}{2}\eta^{2}\right)$
$\propto\exp\left(-\frac{1}{2}\left[(n_{1}+n_{2}+\delta_{1})\phi^{2}+(n_{2}+\delta_{2})\eta^{2}+2n_{2}\phi\eta-2\phi(S_{z}+S_{w})-2\eta S_{w}\right]\right).$
Thus $(\phi,\eta)^{\top}$ follows a bivariate normal distribution: $(\phi,\eta)^{\top}\,|\,z,w\sim N({\mu},{\Sigma})$ where the covariance and mean vector are given as follows:
${\Sigma}=\frac{1}{D}\begin{pmatrix}n_{2}+\delta_{2}&-n_{2}\\ -n_{2}&n_{1}+n_{2}+\delta_{1}\end{pmatrix},\quad D=(n_{1}+n_{2}+\delta_{1})(n_{2}+\delta_{2})-n_{2}^{2}.$
The posterior mean vector is given by
${\mu}={\Sigma}\begin{pmatrix}S_{z}+S_{w}\\ S_{w}\end{pmatrix}=\frac{1}{D}\begin{pmatrix}(n_{2}+\delta_{2})(S_{z}+S_{w})-n_{2}S_{w}\\ -n_{2}(S_{z}+S_{w})+(n_{1}+n_{2}+\delta_{1})S_{w}\end{pmatrix}$
Explicitly,
$\mu_{\phi}=\frac{(n_{2}+\delta_{2})S_{z}+\delta_{2}S_{w}}{(n_{1}+n_{2}+\delta_{1})(n_{2}+\delta_{2})-n_{2}^{2}},\quad\mu_{\eta}=\frac{(n_{1}+\delta_{1})S_{w}-n_{2}S_{z}}{(n_{1}+n_{2}+\delta_{1})(n_{2}+\delta_{2})-n_{2}^{2}},$
and
$\operatorname{Var}_{\phi}=\frac{n_{2}+\delta_{2}}{(n_{1}+n_{2}+\delta_{1})(n_{2}+\delta_{2})-n_{2}^{2}},\quad\operatorname{Var}_{\eta}=\frac{n_{1}+n_{2}+\delta_{1}}{(n_{1}+n_{2}+\delta_{1})(n_{2}+\delta_{2})-n_{2}^{2}}.$
True cut posterior
Given the model settings, the cut-posterior for $\phi$ is defined as $p_{\text{cut}}(\phi\,|\,z,w)=p(\phi\,|\,z)$, where ${\phi}\,|\,z$:
${\phi}\,|\,z\sim N\left(\frac{S_{z}}{n_{1}+\delta_{1}},\;\frac{1}{n_{1}+\delta_{1}}\right).$
Conditional on $\phi$, the downstream posterior of $\eta$ is
$\eta\,|\,w,\phi\sim N\left(\frac{S_{w}-n_{2}\phi}{n_{2}+\delta_{2}},\;\frac{1}{n_{2}+\delta_{2}}\right).$
Marginalizing over $\eta\,|\,z$:
$\mathbb{E}[\eta\,|\,w]=\mathbb{E}[\mathbb{E}[\eta\,|\,\phi,w]]=\frac{(n_{1}+\delta_{1})S_{w}-n_{2}S_{z}}{(n_{1}+\delta_{1})(n_{2}+\delta_{2})}.$
The variance of the marginal posterior for $\eta$ is
$\operatorname{Var}({\eta}\,|\,w)$ $=\mathbb{E}\left[\operatorname{Var}({\eta}\,|\,w,{\phi})\right]+\operatorname{Var}\left[\mathbb{E}({\eta}\,|\,w,{\phi})\right]$
$=\frac{1}{n_{2}+\delta_{2}}+\left(\frac{n_{2}}{n_{2}+\delta_{2}}\right)^{2}\,\frac{1}{n_{1}+\delta_{1}}$
$=\frac{1}{n_{2}+\delta_{2}}+\frac{n_{2}^{2}}{(n_{2}+\delta_{2})^{2}(n_{1}+\delta_{1})}.$
Appendix E Estimating Cause-Specific Mortality Fractions in Mozambique for Child population
Figure 6 presents the estimated densities of the cause-specific mortality fractions (CSMFs) for the child population, consistent with those obtained for the neonate population. The sequential Bayes posteriors differ markedly from the corresponding cut-posteriors. Among the cut-Bayes approaches, NeVI-Cut exhibits close alignment with nested MCMC, whereas the parametric cut, based on a Dirichlet variational family, shows limited flexibility and induces bias in estimation.
Figure 6: Posterior distributions of calibrated cause-specific mortality fractions (CSMFs) for children aged 1-59 months, based on VA-only data collected by the Countrywide Mortality Surveillance for Action (COMSA) program in Mozambique. NeVI-Cut effectively captures complex distributional characteristics, such as multimodality and skewness, achieving performance comparable to nested MCMC.